### v1

# Personalize to generalize: Towards a universal medical multi-modality generalization through personalization

Abstract
Personalized medicine is a groundbreaking healthcare framework for the $21^{st}$ century, tailoring medical treatments to individuals based on unique clinical characteristics, including diverse medical imaging modalities. Given the significant differences among these modalities due to distinct underlying imaging principles, generalization in multi-modal medical image tasks becomes substantially challenging. Previous methods addressing multi-modal generalization rarely consider personalization, primarily focusing on common anatomical information. This paper aims to bridge multi-modal generalization with the concept of personalized medicine. Specifically, we propose a novel approach to derive a tractable form of the underlying personalized invariant representation $\mathbb{X}_{h}$ by leveraging individual-level constraints and a learnable biological prior. We demonstrate the feasibility and benefits of learning a personalized $\mathbb{X}_{h}$, showing that this representation is highly generalizable and transferable across various multi-modal medical tasks. Our method is rigorously validated on medical imaging modalities emphasizing both physical structure and functional information, encompassing a range of tasks that require generalization. Extensive experimental results consistently show that our approach significantly improves performance across diverse scenarios, confirming its effectiveness.
1 Introduction
Figure 1: A diagram of medical modalities and individual differences. Individual variations may be significant and warrant further research attention from the medical intelligence community.
Personalized medicine represents a transformative framework for $21^{st}$ century healthcare, tailoring medical treatments to each patient’s unique characteristics [45, 24, 4]. This approach necessitates diverse information, including clinical data such as radiological images. Three-dimensional medical images, generated through specialized techniques and radiopharmaceuticals, excel at highlighting specific anatomical features. Collectively, different medical image modalities provide a comprehensive view of a patient’s structural and functional characteristics. However, this distinction between medical modalities creates significant generalization challenges in medical image analysis, especially when certain modalities may be inaccessible due to an individual’s financial constraints or physical limitations, thereby complicating the effectiveness of personalized medicine.
As illustrated in Fig. 1, contemporary research in medical intelligence is mainly concentrated on structural modalities that depict physical anatomy. This includes Magnetic Resonance Imaging (MRI) scans [53], which use strong magnetic fields and radiofrequency currents yielding distinct sequences, and Computed Tomography (CT) scans [33, 51], which employ X-rays to measure its attenuation. Other studies [49] focus on the functional modalities associated with biochemistry, such as Positron Emission Tomography (PET) scans. PET scans are expensive functional imaging scans that employ radiotracers emitting gamma rays to visualize and measure metabolic processes. The differing imaging principles result in substantial modality gaps, presenting a critical challenge for model generalization. For clarity, we categorize these modalities for generalization tasks into two types: homogeneous generalization, which pertains to generalizing within structural or functional modalities (e.g., T1, T2, T1ce, and Flair in MRI, as shown in Fig. 1); and heterogeneous generalization, which involves generalizing across both structural and functional modalities, such as CT and PET.
As each person is fundamentally different from the average of the population [45], the concept of multi-modality generalization needs comprehensive discussion within the scope of personalization, an area scarcely addressed in previous research. An ideally well-generalized personalized medical model across modalities should (1) provide additional insights derived from all modalities when only a subset is accessible and (2) seamlessly transfer across domains while maintaining the capacity to achieve (1). While pre-training can significantly enhance downstream generalization, some recent approaches concentrate on learning common physical anatomy invariance at the class level [23], which may overlook the individual variations. Another line of research [40, 46, 23] primarily addresses the transferability of single-modal tasks and may not be suitable for multi-modal scenarios. In addition, most research on homogeneous generalization for medical tasks focuses on structural sequences of MRI or CT, employing strategies such as cross-modality transfer [30, 25, 51] or targeting challenges like missing modality segmentation [31, 6, 36, 37, 51]. Heterogeneous generalization presents a greater challenge due to the disparities between structural and functional information. Despite its significance, very few efforts [34] have addressed heterogeneous generalization, mainly focusing on one-directional modality transfer (e.g., PET to CT or MRI), and rarely exploring the model’s generalizability and transferability for other tasks in this context.
To address multi-modality generalization issue for personalization under both homogeneous and heterogeneous settings, we formally introduce the concept of the personalized invariant representation for multi-modal generalization, denoted as $\mathbb{X}_{h}$, and its constraints as outlined in Hypothesis 3.1. Furthermore, personalized invariant $\mathbb{X}_{h}$, which learns aggregated biological information from all possible modalities specific to the individual, is likely to enhance performance across various medical tasks for that person. Building on this hypothesis, this paper proposes a general approach aimed at enhancing the generalization of various medical imaging tasks through personalization. Specifically, our method constructs an approximation of $\mathbb{X}_{h}$ using the learnable biological prior knowledge $\mathbb{O}$, via decomposition, invariance, and equivariance constraints during pre-training (refer to Sec. 3.2). The learned approximation of $\mathbb{X}_{h}$ can then be utilized to enhance performance in downstream generalization tasks, irrespective of whether a domain gap exists between the pre-training data and downstream data.
Importantly, this paper demonstrates that obtaining a personalized invariant representation, $\mathbb{X}_{h}$, is feasible through our approach, and such invariance leads to generalization improvements across various medical tasks. To validate our methodology, we conduct experiments on modality transfer and missing modality segmentation tasks, addressing not only the homogeneous generalization of MRI but also the rarely explored heterogeneous generalization, such as generalization between PET and CT. Our findings reveal that our approach successfully captures comprehensive, personalized information even when only partial modalities are available for a given individual (see Fig. 3). Moreover, extensive experiments on both homogeneous (Sec. 4) and heterogeneous (Secs. 5 and 6) generalization demonstrate that our approach can be adapted for downstream tasks and surpasses current state-of-the-art (SOTA) methods in multiple tasks, validating its superiority. We will publicly release our code, checkpoints, and data splits upon acceptance.
2 Related work
Medical generalization tasks currently concentrates on homogeneous generalization, introducing tasks, such as modality transfer and missing modality segmentation for structural modalities — Flair, T1, T2, and T1ce of MRI- in brain tumor segmentation [53], or between MRI and CT [51] for modality transfer. [34] propose an approach for heterogeneous generalization in terms of modality transfer, but only tailored for transferring PET to CT. In terms of tasks, most current studies focus on one specific generalization task, either segmentation [5, 43, 12, 52, 44, 52] or modality transfer [22, 54, 13, 35, 26, 30, 38, 11, 33, 25, 47, 51]. This paper aims to develop an approach that is feasible for different downstream tasks under both homogeneous and heterogeneous modality generalization. Our approach aims to learn the $\mathbb{X}_{h}$ through pre-training; we list related medical pre-training work [40, 46, 8, 23] here. A notable work among them is [23], which extracts class-specific anatomical invariance. However, they only focus on a single modality. Such single-modality approaches may not be able to construct $\mathbb{X}_{h}$ for improving the generalization across modalities. The generalization in medical images also connects to alignment in multi-domain generalization for natural images [14, 28, 27, 20, 39]. Please refer to Appendix A for a detailed literature review.
3 Learning personalized invariant representation for medical generalization
Preliminaries. In this paper, we denote the encoder as $\mathcal{E}$ and its corresponding decoder as $\mathcal{D}$. For an individual human being $h\in\mathcal{H}$, the corresponding medical images are represented as $X_{h}={X^{i}_{h},X^{j}_{h},\dots,X^{k}_{h}}$, where $i,j,\dots,k\in\mathcal{M}$, and $\mathcal{M}$ represents the set of all possible modality combinations. We denote the intermediate features produced by $\mathcal{E}(X_{h})$ and $\mathcal{E}(X^{i}_{h})$ as $x_{h}$ and $x^{i}_{h}$, respectively. The final layer features from the encoder are represented as $z_{h}$ and $z^{i}_{h}$. The learned approximation of $\mathbb{X}_{h}$ is denoted as ${\mathbb{X}_{h}}^{\prime}$. Additionally, we define the geometric warping function $\phi^{i}\in\Phi$, where $\phi^{i}(X^{i}_{h})\in\mathcal{X}$, and $\Phi$ denotes the set of all possible geometric warping functions. Finally, $I(\cdot;\cdot)$ and $P(\cdot)$ represent mutual information and probability distribution, respectively.
Before addressing the problem for both homogeneous and heterogeneous generalization, we introduce the $\mathbb{X}_{h}$ Hypothesis for medical imaging:
Hypothesis 3.1 ( $\mathbb{X}_{h}$ Hypothesis).
Consider the set $\mathcal{M}$ of all possible modality combinations and the set $\Phi$ of all possible geometric transformations (e.g., SO(3)) transformations, for example, rotations corresponding to different poses of the person. There exists a personalized invariant representation $\mathbb{X}_{h}$ for an individual from the population $h\in\mathcal{H}$, which can be decomposed into modality-specific images $X^{i}_{h}$ given a modality combination $i\in\mathcal{M}$ as a condition:
$X^{i}_{h}=\mathbb{X}_{h}|i;\;i\in\mathcal{M},h\in\mathcal{H},\quad\textit{s.t.,}\mathbb{X}_{h}\perp\!\!\!\perp\mathcal{M},\Phi.$ (1)
Despite potential differences in modalities and individual variations, clinical diagnoses focus on the biological conditions of a certain patient, which remain mostly invariant during a single hospital visit. Thus, the $\mathbb{X}_{h}$ Hypothesis holds in most cases. Our method aims to obtain an accurate approximation of $\mathbb{X}_{h}$. The overall learning framework for $\mathbb{X}_{h}$ is illustrated on the left-hand side of Fig. 1. Data from each modality are encoded by $\mathcal{E}$, and the encoded features are used to retrieve knowledge from the learnable biological prior $\mathbb{O}$. The features and retrieved knowledge are then fused. By applying constraints of decomposition, equivariance, and invariance on the fused features, we approximate $\mathbb{X}_{h}$ effectively.
Figure 2: Left: Overall framework of learning $\mathbb{X}_{h}$. Right: Diagrams of differences between previous methods of learning $\mathbb{Z}_{h}$ and our proposed method of learning $\mathbb{X}_{h}$.
As illustrated in Fig. 2 right-hand side top, previous approaches [31, 6, 36, 37] learn invariant representations $\mathbb{Z}_{h}$ across modalities through the encoder $\mathcal{E}$ for generalization: $\mathcal{E}(X^{m}_{h})\to\mathbb{Z}_{h},\mathbb{Z}_{h}\perp\!\!\!\perp\mathcal{M},m\in\mathcal{M},h\in\mathcal{H}$ during pre-training or training. Is $\mathbb{Z}_{h}$ a good approximation of $\mathbb{X}_{h}$, and does it benefit the generalization of different downstream tasks? The answer might be negative because such an approach may erase modal-specific information in $\mathbb{Z}$, making it impossible to be decomposed back into different modalities as shown in Eq. 1. Moreover, while current studies [19, 41, 52, 12] also disentangle modality-dependent features alongside the invariant representation $\mathbb{Z}$ to enhance transferability, this strategy may compromise the generalization ability of $\mathbb{Z}$. The reason is that the transferred targets become constrained by the learned modal-dependent features, potentially limiting their broader applicability.
3.1 Using global prior for better personalized invariant representation
To learn a better approximation of $\mathbb{X}_{h}$, we leverage a global biological prior, denoted as $\mathbb{O}$. If $\mathbb{O}$ can be learned, representations from any modality can complete themselves by retrieving the missing knowledge from $\mathbb{O}$, forming a better approximation of $\mathbb{X}_{h}$. Empirically, we initialize a learnable tensor as $\mathbb{O}$. As shown in Fig. 1, the representation $z^{i}_{h}$ retrieves its missing knowledge from $\mathbb{O}$ via attention: ${z^{i}_{h}}^{\prime}:=attn(query:z^{i}_{h},key:\mathbb{O},value:\mathbb{O})$. The original representation and the retrieved knowledge are then fused through convolution: $\mathbb{X}^{i}_{h}:=conv({z^{i}_{h}}^{\prime},z^{i}_{h})$. If the model is well-trained under the constraints of equivariance, invariance, and decomposition, the fused feature $\mathbb{X}^{i}_{h}$ becomes $\mathbb{X}^{\prime}_{h}$, a good approximation of $\mathbb{X}_{h}$. The details of these constraints are discussed in Sec. 3.2.
3.2 Constraints of equivariance, invariance, and decomposition
Contrastive learning. Before we introduce the constraints, we include the contrastive loss as our baseline. During the pre-training stage, we follow previous work [8, 40] and employ the contrastive learning loss. Specifically, the positive pairs are constructed as augmented samples from the same sub-volume, while the negative pairs are the views from different sub-volumes. Similar to [40], the contrastive coding is obtained by attaching a linear layer $\psi(\cdot)$ to the $z_{h}$, its positive pair $z^{+}_{h}$, and all samples $\{z_{h}^{i}\}_{i=1}^{B}$ where $B$ is the total number of samples. The contrastive loss is then defined as:
$\mathcal{L}_{contr}=-\log\frac{\exp\left(\operatorname{sim}\left(\psi(z_{h}),\psi(z^{+}_{h})\right)/t\right)}{\sum_{i=0}^{B}\exp\left(\operatorname{sim}\left(\psi(z_{h}),\psi(z^{i}_{h})\right)/t\right)},$ (2)
where $t$ is the measurement of the normalized temperature scale and $sim(\cdot,\cdot)$ denotes the dot product between normalized embeddings as the similarity.
As discussed in Sec. 3.1, the $\mathbb{X}_{h}$ can be obtained through a model trained under the constraints of equivariance, invariance, and decomposition. The following part presents details of those constraints according to the $\mathbb{X}_{h}$ hypothesis.
Invariance constraint. We constrain the invariance for $\mathbb{X}_{h}$ where $\mathbb{X}_{h}\perp\!\!\!\perp\mathcal{M},\Phi$ through alignment. The $z^{i}_{h}$ firstly uses attention to fetch the knowledge from the prior: ${z^{i}_{h}}^{\prime}=attn(z^{i}_{h},\mathbb{O})$ and then they are concatenated and fused through convolution ${\mathbb{X}^{i}_{h}}^{\prime}=conv(z^{i}_{h}\oplus{z^{i}_{h}}^{\prime})$. Despite the different modality combinations and geometric transformations, $\mathbb{X}_{h}$ should be invariant for the person:
$\mathcal{L}_{inv}=\sum\nolimits\left\|{\mathbb{X}^{i}_{h}}^{\prime},{\mathbb{X}_{h}}^{\prime}\right\|^{2},\quad i\in\mathcal{M}.$ (3)
While it is well aligned, ${\mathbb{X}^{i}_{h}}^{\prime}={\mathbb{X}^{j}_{h}}^{\prime}=...={\mathbb{X}_{h}}^{\prime}$ where $j\in\mathcal{M}$. Empirically, we use ${\mathbb{X}_{h}}^{\prime}\triangleq mean({\mathbb{X}^{i}_{h}}^{\prime},{\mathbb{X}^{j}_{h}}^{\prime},...)$ and $mean(\cdot)$ refers the averaging of the input sequence.
Equivariance constraint. To learn better $\mathbb{O}$ and ${\mathbb{X}_{h}}^{\prime}$ as the personalized invariant representation, we constrain the geometric equivariance and representation invariance. Consider the sample space of all modalities $X^{i}_{h}\in\mathcal{X},i\in\mathcal{M}$, the geometric equivariance constraint forces that the geometry of the generated medical image is equivariant to $\phi^{i}$, which can be constrained by the MSE loss in Eq. 6. Furthermore, such equivariance demands that $\phi(x^{i}_{h})$ and $z^{i}_{h}$ contain the information of the geometric transformation $\phi^{i}$, inferring that it is able to extract the $\phi^{i}$ from $\phi^{i}(x^{i}_{h})$ and $z^{i}_{h}$. Therefre, if $\phi^{i}$ can be extracted from the last-layer output $z^{i}_{h}$, it can also be extracted from the $\phi^{i}(x^{i}_{h})$ from the previous layers:
$\min_{\mathcal{D},\mathcal{F}}Dis(\phi^{i},\mathcal{F}(z^{i}_{h})),$ (4)
where $\mathcal{F}:\mathcal{F}(z^{i}_{h})\to{\phi^{i}}^{\prime}$ extracts the geometric transformation and $Dis(\dot{,}\dot{)}$ denotes the distance measurement between ${\phi^{i}}^{\prime}$ and $\phi^{i}$. Empirically, following [40], we also adopt rotation as the geometric transformation, predicting the angle categories of input sub-volume is rotated. Under this case, $\Phi_{R}$ is defined as rotations at $[0,90,180,270]$ degrees along the z-axis, and $\phi_{r}^{i}\in\Phi_{R}$ is the ground truth rotation categories. $\mathcal{F}{z^{i}_{h}}$ produces the softmax probabilities of rotation categories. The loss is in the form of:
$\mathcal{L}_{equ}=-\sum\nolimits_{r=1}^{|\Phi_{R}|}\phi_{r}^{i}\log\mathcal{F}(z^{i}_{h}).$ (5)
Task T1→T2 T2 → Flair T1 → T1ce
Method PSNR↑ NMSE↓ SSIM↑ PSNR↑ NMSE↓ SSIM↑ PSNR↑ SSIM↑
2D
Pix2Pix [22] 24.624 0.109 0.874 24.82 0.0250 0.846 27.05 0.858
CycleGAN [54] 23.535 0.155 0.837 23.418 0.164 0.825 30.13 0.906
NICEGAN [7] 23.721 0.148 0.840 23.643 0.148 0.829 - -
GcGAN [13] - - - 29.98 - 0.917 25.98 0.872
CUT [35] - - - 23.54 - 0.819 26.27 0.846
RegGAN [26] 24.884 0.094 0.881 24.576 0.112 0.852 31.36 0.930
ResViT [10] 25.578 0.088 0.895 24.825 0.108 0.861 31.46 0.932
Diffusion [11] - - - 31.98 - 0.930 29.22 0.921
MD-Diff [47] - - - 30.76 - 0.934 33.08 0.948
3D
Pix2Pix 23.740 0.138 0.835 23.508 0.152 0.822 - -
CycleGAN 25.181 0.097 0.887 24.602 0.113 0.854 - -
EaGAN [50] 24.884 0.094 0.881 24.576 0.112 0.852 - -
MS-SPADE [25] 25.818 0.079 0.904 25.074 0.098 0.867 26.119 0.912
Ours 30.756 0.065 0.944 32.224 0.046 0.941 34.547 0.955
Table 1: Modality transfer results of MRI on BRATS23: Comparison between previous methods and our method. The best results are highlighted in blue. Results denoted with are gained from [25], while the results denoted with are gathered from [47].
Target T1 T1ce T2 Flair
Source PSNR↑ NMSE↓ SSIM↑ PSNR↑ NMSE↓ SSIM↑ PSNR↑ NMSE↓ SSIM↑ PSNR↑ NMSE↓ SSIM↑
SwinUNETR 32.815 0.092 0.941 31.655 0.202 0.912 24.650 0.361 0.857 27.593 0.202 0.883
MS-SPADE 29.001 0.055 0.942 26.119 0.078 0.912 25.818 0.103 0.904 24.842 0.113 0.859
T1 Ours 43.472 0.003 0.996 34.547 0.045 0.955 30.756 0.065 0.944 31.693 0.049 0.937
SwinUNETR 32.456 0.100 0.929 33.001 0.156 0.926 25.125 0.366 0.859 27.699 0.211 0.882
MS-SPADE 26.228 0.076 0.922 28.759 0.060 0.937 25.990 0.092 0.907 25.204 0.092 0.881
T1ce Ours 34.077 0.020 0.962 46.663 0.003 0.996 30.775 0.063 0.942 32.224 0.046 0.941
SwinUNETR 30.102 0.171 0.896 30.354 0.283 0.883 26.831 0.268 0.887 27.234 0.242 0.872
MS-SPADE 25.422 0.085 0.908 25.234 0.087 0.895 29.230 0.048 0.942 25.074 0.098 0.867
T2 Ours 32.646 0.028 0.955 33.857 0.051 0.949 43.653 0.006 0.991 32.224 0.046 0.941
SwinUNETR 31.371 0.135 0.916 31.285 0.240 0.905 25.579 0.338 0.867 29.092 0.148 0.923
MS-SPADE 25.186 0.090 0.905 25.899 0.094 0.906 26.146 0.086 0.913 28.608 0.058 0.938
Flair Ours 32.752 0.026 0.951 33.471 0.055 0.944 30.571 0.068 0.940 43.624 0.004 0.995
Table 2: Modality transfer results of MRI on BRATS23: Comparison between the previous method and ours for transfer between all four modalities. The averaged results of metrics for all validation samples are listed. The best results are highlighted in blue. Please refer to Appendix 9 for the standard derivations according to each result.
Decomposition constraint. As shown in Eq. 1 of $\mathbb{X}_{h}$ Hypothesis, the ${\mathbb{X}_{h}}^{\prime}$ need to be able to be decomposed as different modalities, which refers: $\min_{\mathcal{E},\mathcal{D},\mathbb{O}}I(P({\mathbb{X}_{h}}^{\prime}|i);P(X^{i}_{h}))$. An intuitive approach is reconstructing all possible modalities by using $\mathbb{X}_{h}$, whose objective can be formed as:
$\mathcal{L}_{decom}\!=\!\sum_{1}^{|\mathcal{M}|}\left\|{\phi^{i}}^{-1}\!\left(\mathcal{D}\left({\mathbb{X}_{h}}^{\prime}|\phi^{i}(x^{i}_{h})\right)\right),X_{h}\right\|^{2},$ (6)
where $i\in\mathcal{M}$ and $\phi^{i}(x^{i}_{h})$ represents intermediate representations produced during $\mathcal{E}(\phi^{i}(X^{i}_{h}))$ and $X_{h}$ denotes all possible modalities. Intuitively, $\phi^{i}(x^{i}_{h})$ from earlier layers of the encoder constrains modality information thus $\mathcal{D}({\mathbb{X}_{h}}^{\prime}|\phi^{i}(x^{i}_{h}))\triangleq\mathcal{D}({\mathbb{X}_{h}}^{\prime}|i)$. Specifically, the generated medical image is transformed back by using the inverse of $\phi^{i}$ to align with the inputs.
Final loss for learning $\mathbb{X}_{h}$. The final loss for pre-training is the combination of the above losses:
$\mathcal{L}_{pre}=\mathcal{L}_{contr}+\mathcal{L}_{decom}+\mathcal{L}_{equ}+\mathcal{L}_{inv},$ (7)
where the weight of each loss is omitted here.
3.2.1 The connection between the constraints and global biological prior
It is important to note that the above constraints are closely interconnected, as they align with Eq. 1. After obtaining additional knowledge from $\mathbb{O}$, the invariance constraint ensures that the representations from each modality for a given individual are the same, such that ${\mathbb{X}^{i}_{h}}^{\prime}$ and ${\mathbb{X}_{h}}^{\prime}$ can be considered equivalent. Combined with the decomposition constraint, which enforces that ${\mathbb{X}_{h}}^{\prime}$ is shared for the generation of all possible modalities, ${\mathbb{X}_{h}}^{\prime}$ is thus able to generalize across modalities.
Additionally, the equivariance and decomposition constraints implicitly maintain SO(3)-equivariance by satisfying the relation $\mathcal{D}\circ\mathcal{E}(\phi^{i}(X^{i}_{h});\theta)=\phi^{i}(\mathcal{D}\circ\mathcal{E}(X^{i}_{h};\theta))$ [^1] The SO(3) transformations are left-multiplication; they are expressed here in a simplified form, using $\phi^{i}(\cdot)$., where $\theta$ represents the model parameters after training. This ensures that geometric transformations are preserved in the latent features $z^{i}_{h}$, such that $\mathcal{F}(z^{i}_{h})=\phi^{i}$. The invariance constraint then requires that $z^{i}_{h}$ with geometric transformations can retrieve features from $\mathbb{O}$ to form ${\mathbb{X}^{i}_{h}}^{\prime}$, which remains invariant to any geometric transformation. This implicitly constrains $\mathbb{O}$ to contain comprehensive biological information, including other potential geometric transformations, thereby improving the $\mathbb{X}_{h}^{\prime}$ through $\mathbb{O}$ in approximating $\mathbb{X}_{h}$ and enhancing the robustness of ${\mathbb{X}_{h}}^{\prime}$.
3.3 Application of different modalities and tasks
After pre-training with the loss function $\mathcal{L}_{pre}$, the model is then utilized for downstream tasks such as segmentation or generation. We denote the commonly used loss functions for these tasks, such as dice loss, cross-entropy loss, or mean squared error (MSE) loss, as $\mathcal{L}_{ori}$, where paired data and labels $(X,Y)\in(\mathcal{X},\mathcal{Y})$ are provided. In addition to $\mathcal{L}_{ori}$, we incorporate the invariance loss, denoted as $\mathcal{L}{inv}$, as part of the fine-tuning process for downstream tasks:
$\mathcal{L}_{down}=\mathcal{L}_{ori}+\mathcal{L}_{inv}.$ (8)
Empirically, we adopt the SwinUNETR architecture [17] as the backbone of the encoder $\mathcal{E}$, and implement the proposed components. The model is trained with $\mathcal{L}_{pre}$ during the pre-training phase; users have the option to either use the standard SwinUNETR by loading only our pre-trained encoder weights or to employ our proposed model structure with all pre-trained weights for downstream tasks. Notably, all modalities for a given individual, $X_{h}={X^{i}_{h},X^{j}_{h},...,X^{k}_{h}},i,j,...,k\in\mathcal{M}$, share the same encoder, with the encoder’s channel size set to match the number of modality types. The input volume size for all experiments is fixed at $96\times 96\times 96$. Further empirical details on how $\mathbb{X}_{h}$ is leveraged for homogeneous and heterogeneous generalization are provided in Secs. 4 and 5.
4 Homogeneous generalization: MRI
Figure 3: Visualization of the efficacy of prior $\mathbb{O}$. Displayed are the generated modalities on the input Flair modality of a testing sample on the BTATS21 dataset. Columns show: the generated images of the model (1) without prior $\mathbb{O}$ and (2) with prior $\mathbb{O}$ are aligned with (3) the GT images. Typically, the differences between without and with prior $\mathbb{O}$ (the (2)-(1) column) are visualized to compare with the differences between without $\mathbb{O}$ and GT (the (3)-(1) column). Red and blue refer to the positive (accomplishment) and negative (refinement) values of the differences, respectively.
This section demonstrates that our approach enhances the homogeneous generalization across structural modalities in MRI. To validate that our method captures personalized information, especially anatomical structure features, during the pre-training stage, we first apply it to modality transfer tasks. Next, we adapt the pre-trained model for the downstream missing modality segmentation task. Experimental results indicate that our approach outperforms state-of-the-art (SOTA) methods in both tasks, thereby supporting the $\mathbb{X}_{h}$ Hypothesis and confirming the effectiveness of our method.
4.1 Pre-training for modality transfer
Modality transfer tasks focus on converting medical images from multiple modalities to other modalities. We test our approach on the structural modalities of MRI. This task aligns seamlessly with our pre-training objective, and the quality of the generated modalities serves as a validation of the anatomical knowledge captured by our pre-trained model. Importantly, our approach can generate all modalities without knowing the exact modalities where the input form, as the learned representation ${\mathbb{X}_{h}}^{\prime}$ encompasses comprehensive information across all possible modalities.
All settings FM MN=1 MN=2 MN=3
Method Mean Std. - Mean Std. Mean Std. Mean Std.
Tumor core
RFNET [12] 76.08 6.99 83.40 80.63 4.53 76.57 7.15 68.95 6.07
mmFormer [52] 76.43 5.83 82.22 79.78 4.33 76.55 6.03 71.45 6.80
SPA [42] 74.80 6.95 82.23 78.99 5.38 75.01 8.31 68.44 8.88
M3AE [29] 72.67 7.43 80.29 77.61 4.59 73.37 7.96 64.79 9.85
M2F [38] 73.69 6.83 80.34 77.48 5.19 74.17 6.88 67.51 6.63
Ours 79.78 4.55 86.72 83.64 2.29 79.56 3.28 74.51 3.87
Enhancing tumor
RFNET 59.31 15.10 73.65 66.91 12.90 59.17 14.50 48.35 13.86
mmFormer 62.14 18.60 79.91 71.54 15.94 61.77 19.65 48.86 20.73
SPA 58.92 17.68 73.40 68.44 17.11 58.05 19.77 47.10 19.99
M3AE 55.98 17.45 73.79 65.09 14.54 55.53 20.37 43.09 22.03
M2F 58.84 16.58 75.26 66.67 13.83 58.99 16.19 46.70 15.97
Ours 63.49 7.58 70.64 64.44 6.62 63.87 11.45 60.19 11.31
Whole tumor
RFNET 83.92 6.14 89.27 87.25 2.94 84.95 3.81 77.70 7.46
mmFormer 84.84 5.35 88.26 87.59 2.40 85.36 5.60 80.45 4.64
SPA 84.52 5.48 89.03 87.81 1.25 85.26 4.69 78.98 7.32
M3AE 81.52 6.71 86.82 85.64 1.36 82.43 6.08 74.74 8.69
M2F 83.88 5.79 88.72 87.30 1.99 84.62 2.57 78.13 7.81
Ours 87.63 3.25 91.19 89.49 1.82 87.45 1.03 85.17 3.93
Table 3: Missing modality segmentation results of MRI on BRATS18: Num denotes the number of missing modalities for different settings. We report mean and standard deviations of DICE results of all experimental combinations under the same Num. The best results are highlighted in blue. Please refer to Appendix Tab. 10 for detailed results of each setting. MN: Missing Num; FM: Full Modality.
SSIM↑ PSNR↑
+ Contrastive
+ Decomposition
$\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$
+ Equivariance $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$
+ Invariance $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$
+ $\mathbb{O}$bullet $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$
PET $\to$ PET 0.9903 0.9835 0.9955 0.9931 0.9957 0.9969 44.8811 42.2223 46.5603 45.5829 47.4198 49.5473
CT $\to$ CT 0.9739 0.9475 0.9419 0.9437 0.9664 0.9780 37.2309 32.0777 31.2692 33.1866 35.4194 37.0989
PET $\to$ CT 0.9161 0.9148 0.9215 0.9070 0.9121 0.9282 28.1046 29.3181 29.6694 26.8885 27.2708 30.1548
CT $\to$ PET 0.9884 0.9824 0.9851 0.9834 0.9842  0.9883 39.8490 39.0795 39.1718 39.4528 39.4348 41.5840
Avg. 0.9672 0.9571 0.9610 0.9568 0.9646 0.9728 37.5164 35.6744 36.6677 36.2777 37.3862 39.5963
Table 4: Ablation study - Modality transfer results of PET and CT on AutoPET-II: Ablation results of models trained under different combinations of constraints. The best and second results are highlighted in blue and cyan, respectively.
Experimental settings. Following previous methods [25], we utilize the multi-modal brain tumor segmentation challenge 2023 (BRATS23) dataset [1, 32, 3, 2]. BRATS23 includes four structural MRI modalities (T1, T1ce, T2, and FLAIR) for each individual. Our model is tested on the BRATS23 validation set, which contains these four modalities for 219 individuals. We evaluate the quality of synthesis using peak signal-to-noise ratio (PSNR), normalized mean squared error (NMSE), and structural similarity index (SSIM) [48]. To provide comprehensive results, we separately compare the translation results for T1 $\to$ T2 and T2 $\to$ FLAIR, as some previous methods are only capable of single-modality transfer. These include both 2D and 3D generation methods, as shown in Tab. 1. Additionally, we employed SwinUNTER for multi-modality translation comparisons. All evaluations were performed on 3D volumes; for the 2D methods, synthesized target images were stacked to form a 3D volume for comparison. Please refer to model training details to Appendix D.
Results. Tab. 1 exhibits the transfer results of T1 $\to$ T2 and T2 $\to$ Flair. Our approach significantly surpasses previous 2D and 3D generation methods, including single- and multi-modality translation methods. Specifically, our approach exceeds current SOTA diffusion-based methods, such as 2D-based MD-Diff and 3D-based MS-SPADE. In terms of multi-modality translation, as Tab. 2 shows, our approach performs better than MS-SPADE and SwinUNETR across all metrics under all settings. Moreover, for all settings, it can be seen that our method significantly improves the SSIM, indicating a better anatomy structure obtained by our approach. These results indicate that the $\mathbb{X}_{h}$ Hypothesis is plausible for homogeneous generalization, and our personalized approach is able to obtain its approximation.
Analysis of $\mathbb{O}$. We show that using $\mathbb{O}$ for $\mathbb{X}_{h}$ mainly accomplishes the personalized knowledge of each sequence from MRI modalities. Those modalities are mainly focused on the physical anatomy. For the Flair modality in MRI, which mainly highlights the lesion but suppresses structures like bones, Fig. 3 shows that without $\mathbb{O}$, the main difference between the generated images and ground truth (GT) images is the personalized structure. Prior $\mathbb{O}$ for $\mathbb{X}_{h}$ accomplishes and refines the personal level anatomical information, mitigating the gap between them with the GT, so it can be better transferred to other structural focusing modalities.
Method Dice↑ Dice-↑ TPR↑ TNR↑ FNR↓ FPR↓
From scratch
nnUnet [21] 33.10 - - - - -
UNETR [18] 10.81 23.14 80.65 3.64 19.35 96.36
SwinUNETR [17] 43.45 62.60 90.32 62.73 9.68 37.27
SwinUNETR with pre-train
With pre-train [40] 44.06 57.79 89.25 73.64 10.75 26.36
With ours 48.20 61.16 88.17 77.27 11.83 22.72
Table 5: Segmentation results of PET and CT on AutoPET-II: Comparison between the previous method and ours. The best results are highlighted in blue.
4.2 Tuning for missing modality segmentation
Experimental settings. To validate the generalization ability of the pre-trained model, we fine-tune the model obtained from Sec. 4.1 on the BRATS18 [32] from the Multimodal Brain Tumor Segmentation Challenge. Similar to BRATS23, BRATS18 also consists of the same four structural modalities. We employ the Dice similarity coefficient (DICE) as the metric for evaluation. For a fair comparison, we follow data splits of [38] and reproduce the results of previous methods  on these splits by using their released code and following their original settings  [^2] Though we tried our best, it can be noticed some reproduced results are lowered than their reported results in their original paper. It should be clarified that our results also exceed those reported results. However, for a comprehensive study, we mainly report our reproduced results. . Additional experimental details can be seen in Appendix D.
Results. Tab. 3 presents the segmentation results of our approach compared to previous methods. We also compute the standard deviation of DICE scores under various missing modality settings, which highlights the robustness of our model. Notably, our approach outperforms previous methods in most missing modality scenarios, particularly when the number of missing modalities is large. Our method shows significant improvements in enhancing tumor segmentation, especially when Missing Num = 3. Moreover, the reduced standard deviation of DICE scores under different missing modality settings indicates that our personalized approach consistently delivers superior segmentation results. This performance improvement stems from the enhanced generalization of our model, which is rooted in the learned ${\mathbb{X}_{h}}^{\prime}$.
5 Heterogeneous generalization: PET and CT
Given the differing imaging principles, the modality gap in heterogeneous generalization may be more pronounced than that in homogeneous generalization, making the former tasks more challenging. In this section, we evaluate our approach for heterogeneous generalization.
5.1 Pre-training for modality transfer
Experimental settings. We utilize the AutoPET-II dataset [15] from the Automated Lesion Segmentation in PET/CT challenge for pre-training. The AutoPET-II dataset includes AC-PET and CT pairs, where the PET scans adopt FDG tracers, and their attenuation is corrected using the corresponding CT scans. Specifically, we divide the AutoPET-II dataset into training and testing sets. Similar to our approach for heterogeneous generalization, we adopt the Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) as evaluation metrics. In this section, we present the results of models that employ different combinations of the constraints and the set $\mathbb{O}$. We use both contrastive loss and the decomposition constraint as our baseline. Please refer to training details in Appendix D.
SSIM↑ PSNR↑
HNSCC validation NAC→CT NAC →AC Avg. NAC→CT NAC→AC Avg.
UNETR 0.4899 0.8998 0.6949 21.7330 42.8557 32.2944
SwinUNETR 0.5853 0.9265 0.7559 23.5628 42.5495 33.0561
Ours 0.6939 0.9516 0.8227 25.8498 46.4658 36.1578
SSIM↑ PSNR↑
NSCLC NAC→CT NAC →AC Avg. NAC→CT NAC→AC Avg.
UNETR 0.4476 0.8703 0.6590 20.6182 40.8570 30.7376
SwinUNETR 0.4476 0.8705 0.6591 22.5086 41.3272 31.9179
Ours 0.4744 0.8853 0.6798 22.7791 42.7687 32.7739
Table 6: Modality transfer results of NAC-PET to AC-PET and CT that tuned on HNSCC and evaluated on HNSCC validation set and NSCLC: Comparison between the previous method and ours for transfer between different modalities. The best results are highlighted in blue.
Figure 4: Modality transfer results of NAC-PET to AC-PET: Generated examples on the NSCLC dataset for NAC $\to$ AC across individuals.
Figure 5: Generated samples on AUTOPET-II.
Results. As shown in Tab. 4 and generated examples in Fig. 5, incorporating $\mathbb{O}$ with different combinations of constraints improves generation quality across most metrics. Specifically, using the constraints without $\mathbb{O}$ does not guarantee improvements, as discussed in Sec. 3.2.1. Ultimately, employing all constraints along with $\mathbb{O}$ yields the best average results across all translations, validating that our approach performs well in heterogeneous generalization settings. These results indicate that our method under the scope of personalization bridges the gap between structural and functional modalities. We validate the transferability of all these pre-train models in Sec. C.1, where additionally analysis are provided.
5.2 Tuning for segmentation
Experimental settings. We utilize the AutoPET-II [15] dataset for segmentation, evaluating performance using the DICE metric. It is important to note that we employ the same training and testing splits as in Sec. 5.1 to avoid data leakage. Specifically, we adhere to the settings from the official challenge; DICE is calculated in the standard manner but is set to zero for false negatives and true negatives. Additionally, we introduce DICE- to include the mean across all samples, along with true positive rate (TPR), true negative rate (TNR), false negative rate (FNR), and false positive rate (FPR) for the missing modality segmentation evaluation. Our method is compared against nnUNET [21], UNETR [18], and SwinUNETR [17], which are trained directly on the dataset without pre-training. Notably, we also compare our approach with SwinUNETR using its pre-training strategy [40]. Please refer to training details in Appendix D.
Results. The full modality segmentation results are exhibited in Tab. 5. The results indicate that with proper model architecture, such as SwinUNETR, using both two modalities usually outperforms solely using PET. It can be observed that models using our pre-train improve the results across all metrics. Typically, SwinUNETR using our pre-train significantly exceeds it without our pre-trained model, indicating the personalized invariant learned by our pre-train generalizes to the downstream well and can boost the downstream tasks. Moreover, using our proposed components with the pre-train leads to the best DICE and DICE-. This validates that using the prior further emphasizes the personalized invariant, which yields the most segmentation improvements.
6 Special case: A complex scenario
We introduce a more complex scenario, in which the pre-trained model for heterogeneous generalization settings is tuned downstream that span both heterogeneous and homogeneous generalization.
Experimental settings. The pre-train model we adopted is from Sec. 5 that trained on AC-PET and CT. Specifically, we tune the model by using the Head and Neck Squamous Cell Carcinoma (HNCSS) dataset [16] as the training set and the Non-Small Cell Lung Cancer (NSCLC) dataset as the testing set. Both datasets are sourced from The Cancer Imaging Archive (TCIA) [9], and they contain paired non-attenuation-corrected PET (NAC-PET), attenuation-corrected PET (AC-PET), and CT scans. The model is pre-trained for heterogeneous generalization between AC-PET and CT. It is tuned for both homogeneous generalization between AC-PET and NAC-PET and heterogeneous generalization between NAC-PET and CT. Similar to the previous translation experiments, we use SSIM and PSNR as evaluation metrics. Performance in this scenario further validates the model’s generalization capabilities. Note here the training and testing data in the downstream task come from different domains. See training details in Appendix D.
Results. Tab. 6 presents the results on the HNSCC dataset, while Fig. 4 displays generated sample images for homogeneous generalization. Our approach achieves superior results across both heterogeneous and homogeneous generalizations. For heterogeneous generalization, our method consistently improves SSIM for NAC-PET to CT, indicating that the learned ${\mathbb{X}_{h}}^{\prime}$ successfully captures and emphasizes anatomical structures in the generated images, as indicated by improved SSIM. Moreover, though the model is pre-trained between AC-PET and CT, the improvements are also consistent for NAC-PET and AC-PET. These findings confirm that our personalized approach is effective for a complex real-world scenario, demonstrating the transferability and generalizability of the pre-trained model to downstream tasks under various scenarios.
7 Conclusion
This paper proposes a universal approach to address multi-modality generalization by approximating a personalized invariant representation, $\mathbb{X}_{h}$, through constraints of invariance, equivariance, and decomposition, guided by a learnable biological prior. We demonstrate that learning $\mathbb{X}_{h}$ is both feasible and highly beneficial for enhancing generalization in medical tasks. By discussing limitations, future directions, and social impact in Appendices B and C, our study may point to a promising path for achieving medical generalization through personalization in complex settings.
Appendix A Related work
Medical generalization tasks. Most current work focuses on homogeneous generalization, introducing tasks such as modality transfer and missing modality segmentation. The most commonly employed structural modalities — Flair, T1, T2, and T1ce of MRI — are used for brain tumor segmentation [53], or between MRI and CT [51] for modality transfer. [34] propose an approach for heterogeneous generalization in terms of modality transfer, but only tailored for transferring PET to CT.
Self-supervised medical pre-train models for medical generalization. Our approach aims to learn the $\mathbb{X}_{h}$ through pre-training. We list related medical pre-training work [40, 46, 8, 23] here. A notable work among them is [23], which extracts class-specific anatomical invariance. However, they only focus on a single modality. Such single-modality approaches may not be able to construct $\mathbb{X}_{h}$ for improving the generalization across modalities.
Alignment in multi-domain generalization. The issue of cross-modality generalization is similar to the problem of multi-domain generalization, which aims to extract domain invariant representations [14, 28, 27, 20, 39]. Most of these approaches focus on learning invariance across different domains, which may not fit the scope of personalization.
Generalization for medical translation. Typical modality transfer approaches are based on GAN models [22, 54, 13, 35, 26]. In contrast to these GAN-based approaches, some work adopts transformer models [30, 38], while others, such as [11, 33, 25, 47], explore diffusion-based approaches. The methods such as MedM2G [51] further incorporate textual information for modality transfer. Additionally, UNET-like architectures, which can also be applied to these tasks, are highlighted in [18, 17]. Most current modality transfer research focuses on improving synthesis quality. Our approach, however, demonstrates that full-modality transfer, when accompanied by specific constraints, not only enhances generation but also improves downstream generalization.
Generalization for medical segmentation. There are three main types of approaches to missing modality segmentation. Knowledge distillation-based approaches transfer knowledge from models with complete modality information (teachers) to models with missing modality information (students) [5, 43]. [12, 52] recover missing information by leveraging the multimodal latent feature space. Domain adaptation-based methods aim to reduce the gap between models with complete and incomplete modalities by aligning their domains [44]. One prominent shared latent space method, MmFormer [52], exploits intra- and inter-modality dependencies for feature fusion, which is closely related to our work. Our work reveals that our pre-train model with basic segmentation tuning exceeds these approaches.
Appendix B Limitations, challenges, and future work
To enhance the validation of our approach, we adhere to commonly used settings during the tuning stage. Exploring alternative strategies, such as knowledge distillation, could further improve downstream performance. Our approach requires datasets where all modalities are instance-level matched, which can be a stringent condition and may be unattainable for certain modalities. Future research should explore methods to achieve personalized invariance without relying on instance-level matched datasets. Additionally, we advocate for the availability of more open-source multi-modal medical datasets, particularly for functional modalities, as these are not widely accessible to researchers.
Appendix C Social impact
This work presents an approach to tackle multi-modality generalization through personalization. We hope our work can encourage the community to work towards practical, personalized medical models with border generalization ability.
C.1 Downstream segmentation ablation study
Table 7: Ablation study - Segmentation results of using different pre-train models on AutoPET-II: Comparison between the pre-train models with different settings and ours. The best results are highlighted in blue and cyan.
[…]
1 + Contrastive + Decomposition + Equivariance 40.85 55.79 81.72 69.09 18.28 30.91
2 + Contrastive + Decomposition + Invariance 44.34 48.63 77.42 91.82 22.58 8.18
3 + Contrastive + Decomposition + Equivariance + Invariance 42.42 60.67 89.25 63.64 10.75 36.36
4 + Contrastive + Decomposition + Equivariance + $\mathbb{O}$ 46.31 55.77 83.87 82.73 16.13 17.27
5 + Contrastive + Decomposition + Invariance + $\mathbb{O}$ 44.42 57.80 88.17 74.55 11.83 25.45
6
+ Contrastive + Decomposition + Equivariance + Invariance + $\mathbb{O}$
48.20 61.16 88.17 77.27 11.83 22.72
We demonstrate the effectiveness of our proposed components and discuss the process of learning an anatomy-invariant representation. Experimental results for downstream segmentation tasks and visualizations of the pre-trained models are presented in Tab. 7. All experiments are conducted under consistent settings to ensure a fair comparison.
Using all constraints together with $\mathbb{O}$ yields the best results. Consistent with Sec. 3.2.1, the results indicate that using different constraints alone may not guarantee improvements; however, incorporating all constraints along with $\mathbb{O}$ results in the best outcomes. This validates the plausibility of the $\mathbb{X}_{h}$ Hypothesis and demonstrates that achieving good approximation of it significantly enhances generalization.
Using prior $\mathbb{O}$ with decomposition constraint improves the model performance for different settings. Despite different settings, additionally using $\mathbb{O}$ with decomposition improves the downstream model performance. Combined with the improvements from modality transfer results in Tab. 4, it suggests that $\mathbb{O}$ helps with better obtaining anatomical structure.
The invariance and equivariance constraints can not be applied to the same feature. It needs to be highlighted that invariance and equivariance constraints can not be applied to the same features as they conflict with each other. As shown in task 3, without $\mathbb{O}$, invariance and equivariance constraints are applied to the latent feature simultaneously, leading to a significant performance drop. In comparison, apply equivariance constraint before using $\mathbb{O}$ and applying invariance constraint after using $\mathbb{O}$ yields the best results. This is because the geometrical transformation contained in $z^{i}_{h}$ needs to be accomplished by fetching other possible geometrical transformation information form $\mathbb{O}$ and then fused to be invariant.
Appendix D Experimental details
The model and data loaders are built by using MONAI https://docs.monai.io/en/stable/index.html. Please refer to all the details of the implementation in the code. We present some key implementations below.
D.1 Overall training procedure
We provide a pseudo-code for our approach. The loss calculation for Pre-training procedure is simplified as Algorithm 1 and Downstream tuning as Algorithm 2. It is notable that the empirical procedure is flexible as long as the $\mathbb{O}$ is properly used to construct $\mathbb{X}_{h}^{\prime}$ and those constraints are applied to $\mathbb{X}_{h}^{\prime}$.
Data: $X\in\mathcal{X},epoch$
Initialize learnable $\mathbb{O}$ $\mathcal{E}(\cdot),\mathcal{D}(\cdot)$;
while $i\neq epoch$ do
       $X_{h}^{\prime}\leftarrow None$;
       for $h\in\mathcal{H}$ do
             for $i\in\mathcal{M}$ do
                   $\mathcal{L}_{pre}\leftarrow 0$;
                   $X_{h}^{i}\sim X$, $\phi^{i}\sim\Phi$;
                   ${X_{h}^{i}}^{+},{X_{h}^{i}}^{-}=Augment(\phi^{i}(X_{h}^{i}))$;
                   $({z_{h}^{i}},{x_{h}^{i}}),({z_{h}^{i}}^{-},{x_{h}^{i}}^{-}),({z_{h}^{i}}^{+},{x_{h}^{i}}^{+})\leftarrow\mathcal{E}(\phi^{i}(X_{h}^{i})),\mathcal{E}({X_{h}^{i}}^{-}),\mathcal{E}({X_{h}^{i}}^{+})$;
                   Calculate $\mathcal{L}_{contr}({z_{h}^{i}},{z_{h}^{i}}^{+},{z_{h}^{i}}^{-})$, $\mathcal{L}_{pre}+=\mathcal{L}_{contr}$;
                   $\mathcal{F}(z_{h}^{i})\to{\phi^{i}}^{\prime}$;
                   Calculate $\mathcal{L}_{equ}({\phi^{i}}^{\prime},{\phi^{i}})$, $\mathcal{L}_{pre}+=\mathcal{L}_{equ}$;
                   ${z^{i}_{h}}^{\prime}:=Attn(query:z^{i}_{h},key:\mathbb{O},value:\mathbb{O})$;
                   ${X_{h}^{i}}^{\prime}:=Conv(z^{i}_{h}{}^{\prime},z^{i}_{h})$ ;
                   if $X_{h}^{\prime}$ is not None ;
                    /* For saving memory */
                   then
                         Calculate $\mathcal{L}_{inv}(X_{h}^{i}{}^{\prime},X_{h}^{\prime})$, $\mathcal{L}_{pre}+=\mathcal{L}_{inv}$;
                         $X_{h}^{\prime}:=(X_{h}+X_{h}^{i}{}^{\prime})/2$;
                  else
                         $X_{h}^{\prime}:=X_{h}^{i}{}^{\prime}$ ;
                   end if
                  $X_{h}^{i}{}^{\prime}:=\mathcal{D}(X_{h}^{i}{}^{\prime},x_{h}^{i})$;
                   Calculate $\mathcal{L}_{decom}(\phi^{i}{}^{-1}(X_{h}^{i}{}^{\prime}),X_{h})$, $\mathcal{L}_{pre}+=\mathcal{L}_{decom}$;
             end for
       end for
end while
Algorithm 1 Calculate losses during one step for pre-training
Data: $(X,Y)\in(\mathcal{X},\mathcal{Y}),epoch$
Load pre-trained $\mathbb{O}$ $\mathcal{E}(\cdot),\mathcal{D}(\cdot)$;
while $i\neq epoch$ do
       $X_{h}^{\prime}\leftarrow None$;
       for $h\in\mathcal{H}$ do
             for $i\in\mathcal{M}$ do
                   $\mathcal{L}_{down}\leftarrow 0$;
                   $(X_{h}^{i},Y_{h})\sim X,Y$ ;
                   $({z_{h}^{i}},{x_{h}^{i}})\leftarrow\mathcal{E}({X_{h}^{i}})$;
                   ${z^{i}_{h}}^{\prime}:=Attn(query:z^{i}_{h},key:\mathbb{O},value:\mathbb{O})$;
                   ${X_{h}^{i}}^{\prime}:=Conv(z^{i}_{h}{}^{\prime},z^{i}_{h})$ ;
                   if $X_{h}^{\prime}$ is not None ;
                    /* For saving memory */
                   then
                         Calculate $\mathcal{L}_{inv}(X_{h}^{i}{}^{\prime},X_{h}^{\prime})$, $\mathcal{L}_{down}+=\mathcal{L}_{inv}$;
                         $X_{h}^{\prime}:=(X_{h}+X_{h}^{i}{}^{\prime})/2$;
                  else
                         $X_{h}^{\prime}:=X_{h}^{i}{}^{\prime}$ ;
                   end if
                  $Y_{h}^{\prime}:=\mathcal{D}(X_{h}^{i}{}^{\prime},x_{h}^{i})$;
                   Calculate $\mathcal{L}_{ori}(Y_{h}^{\prime},Y_{h})$, $\mathcal{L}_{down}+=\mathcal{L}_{ori}$;
             end for
       end for
end while
Algorithm 2 Calculate losses during one step for fine-tuning
D.2 Homogeneous generalization: structural modalities in MRI
D.2.1 Pre-training and Modality transfer.
Experimental settings. We use four A100 GPUs for training. The learning rate we used for the modality transfer is set to $0.0002$, and the training epoch is set to $1000$. Both the number of input and out channels is set as $4$.
Training details. For the model, both the input and output channels are set to $4$, corresponding to the four MRI modalities. All modalities are loaded and cropped to a size of $96\times 96\times 96$ simultaneously. Following [25], we also normalize each MRI modality to have zero mean and unit variance. During training, the background is excluded for modal generation. A single modality is repeated four times to create four channels during training to obtain ${\mathbb{X}^{i}_{h}}^{\prime}$. The training loss follows the $\mathcal{L}_{pre}$, whose calculation details during the training phase can be seen in Algorithm 1.
D.2.2 Missing modality segmentation.
We use four A100 GPUs for tuning. The learning rate we used for the modality transfer is set to $0.0002$, and the training epoch is set to $1000$. Both the number of input and out channels is set as $4$.
Training details. Following [38], we also normalize each MRI modality to zero mean and unit variance. For the fine-tuning, we employ Dice loss, the weighted cross-entropy loss that is adopted by [38], and the additional $\mathcal{L}_{inv}$.
D.3 Heterogeneous generalization: PET and CT modalities
D.3.1 Modality transfer
All models are trained using A100 GPUs. Training details. All models are trained under the same situations, using the same data pre-processing transforms.
D.3.2 Downstream segmentation
Training details. All training and fine-tuning experiments use the same losses, while the approaches with our pre-train additionally use $\mathcal{L}_{inv}$ for downstream fine-tuning. Moreover, we also compare the original architecture of SwinUNETR using our pre-trained weights with fully using our architecture and our weights for fine-tuning.
D.4 Special case: Tuning from heterogeneous to homogeneous generalization with domain gap
Training details. For the fine-tuning stage, we use the decoder architecture of SwinUNETR, which is randomly initialized. The training procedure is similar to the above modality transfer experiments, with the primary difference being that the input and output channels are set to two. Additionally, we reproduced the results of UNETR and SwinUNETR for comparison, ensuring that the same loss functions were applied across models.
Appendix E More results
E.1 Modality transfer results on BRATS22:
Tab. 8 and Tab. 9 presents the generation result with standard derivations. The results of our method and SwinUNETR are produced by ourselves, while the rest of the results are gathered from [25]. Generated examples are presented in Figs. 6, 7 and 8.
Table 8: Modality transfer results of MRI on BRATS23: Comparison between previous methods and our method. The best results are highlighted in blue.
Task T1→T2 T2 → Flair
Dimension Method PSNR↑ NMSE↓ SSIM↑ PSNR↑ NMSE↓ SSIM↑
Pix2Pix 24.624 ± 0.962 0.109 ± 0.028 0.874 ± 0.015 24.361 ± 1.061 0.117 ± 0.021 0.846 ± 0.019
CycleGAN 23.535 ± 1.334 0.155 ± 0.035 0.837 ± 0.028 23.418 ± 0.944 0.164 ± 0.033 0.825 ± 0.035
NICEGAN 23.721 ± 1.136 0.148 ± 0.029 0.840 ± 0.029 23.643 ± 1.045 0.148 ± 0.022 0.829 ± 0.033
RegGAN 24.884 ± 0.991 0.094 ± 0.024 0.881 ± 0.017 24.576 ± 1.073 0.112 ± 0.022 0.852 ± 0.028
2D ResViT 25.578 ± 0.812 0.088 ± 0.021 0.895 ± 0.018 24.825 ± 1.030 0.108 ± 0.018 0.861 ± 0.021
CycleGAN 25.181 ± 0.861 0.097 ± 0.031 0.887 ± 0.012 24.602 ± 1.181 0.113 ± 0.021 0.854 ± 0.018
Pix2Pix 23.740 ± 1.198 0.138 ± 0.032 0.835 ± 0.019 23.508 ± 1.301 0.152 ± 0.039 0.822 ± 0.024
EaGAN 24.884 ± 0.991 0.094 ± 0.024 0.881 ± 0.017 24.576 ± 1.073 0.112 ± 0.022 0.852 ± 0.028
MS-SPADE 25.818 ± 0.857 0.079 ± 0.016 0.904 ± 0.012 25.074 ± 1.085 0.098 ± 0.021 0.867 ± 0.018
3D Ours 30.756 ± 1.950 0.065 ± 0.034 0.944 ± 0.031 32.224 ± 2.518 0.046 ± 0.029 0.941 ± 0.025
Table 9: Modality transfer results of MRI on BRATS23: The averaged results with standard derivations of metrics between all modalities.
Target T1 T1ce T2 Flair
Source PSNR↑ NMSE↓ SSIM↑ PSNR↑ NMSE↓ SSIM↑ PSNR↑ NMSE↓ SSIM↑ PSNR↑ NMSE↓ SSIM↑
SwinUNETR 32.815 0.092 0.941 31.655 0.202 0.912 24.650 0.361 0.857 27.593 0.202 0.883
Std. 0.968 0.043 0.049 1.062 0.067 0.052 1.008 0.069 0.077 1.144 0.072 0.050
MS-SPADE 29.001 0.055 0.942 26.119 0.078 0.912 25.818 0.103 0.904 24.842 0.113 0.859
Std. 0.643 0.025 0.022 0.816 0.022 0.015 0.857 0.030 0.014 0.728 0.034 0.019
Ours 43.472 0.003 0.996 34.547 0.045 0.955 30.756 0.065 0.944 31.693 0.049 0.937
T1 Std. 2.495 0.004 0.011 1.956 0.030 0.018 1.950 0.034 0.031 2.287 0.024 0.019
SwinUNETR 32.456 0.100 0.929 33.001 0.156 0.926 25.125 0.366 0.859 27.699 0.211 0.882
Std. 1.018 0.044 0.048 0.889 0.055 0.051 0.964 0.071 0.074 1.129 0.071 0.049
MS-SPADE 26.228 0.076 0.922 28.759 0.060 0.937 25.990 0.092 0.907 25.204 0.092 0.881
Std. 0.794 0.027 0.033 0.885 0.019 0.015 0.859 0.032 0.908 0.811 0.050 0.037
Ours 34.077 0.020 0.962 46.663 0.003 0.996 30.775 0.063 0.942 32.224 0.046 0.941
T1ce Std. 2.484 0.012 0.017 3.240 0.004 0.008 1.812 0.030 0.028 2.518 0.029 0.025
SwinUNETR 30.102 0.171 0.896 30.354 0.283 0.883 26.831 0.268 0.887 27.234 0.242 0.872
Std. 1.405 0.056 0.050 1.249 0.086 0.054 1.144 0.054 0.075 1.154 0.073 0.051
MS-SPADE 25.422 0.085 0.908 25.234 0.087 0.895 29.230 0.048 0.942 25.074 0.098 0.867
Std. 0.852 0.026 0.020 1.152 0.034 0.025 0.720 0.018 0.915 1.085 0.021 0.018
Ours 32.646 0.028 0.955 33.857 0.051 0.949 43.653 0.006 0.991 32.224 0.046 0.941
T2 Std. 2.391 0.028 0.028 1.925 0.040 0.027 3.467 0.024 0.038 2.518 0.029 0.025
SwinUNETR 31.371 0.135 0.916 31.285 0.240 0.905 25.579 0.338 0.867 29.092 0.148 0.923
Std. 1.198 0.051 0.054 1.161 0.077 0.053 0.956 0.064 0.073 0.974 0.055 0.049
MS-SPADE 25.186 0.090 0.905 25.899 0.094 0.906 26.146 0.086 0.913 28.608 0.058 0.938
Std. 0.759 0.028 0.048 1.039 0.025 0.027 0.636 0.028 0.944 0.769 0.025 0.028
Ours 32.752 0.026 0.951 33.471 0.055 0.944 30.571 0.068 0.940 43.624 0.004 0.995
Flair Std. 2.399 0.020 0.022 1.634 0.035 0.021 1.951 0.035 0.034 2.441 0.008 0.013
E.2 Missing modality segmentation results on BRATS18:
We provide detailed segmentation results on BRATS18 as Tab. 10.
Table 10: Missing modality segmentation results of MRI on BRATS18: Num denotes the number of missing modalities for different settings. The used modalities are highlighted with gray boxes and the missing ones remain as blank. The results of each setting are presented accordingly.
Missing Num =3 =2 =1 =0
flair
T1
T1ce
Modality T2
SPA 65.86 65.27 78.26 66.4 72.99 83.23 70.66 81.25 70.66 80.63 83.22 73.89 83.36 82.05 83.4
M3AE 69.4 65.45 79.12 71.84 79.9 70.45 82.79 81.17 71.62 73.35 81.78 82.42 73.31 81.61 82.22
mmFormer 67.8 77.32 64.56 64.08 81.51 79.43 69.14 70.63 68.6 80.75 81.75 70.92 81.74 81.55 82.23
RFNET 64.03 74.53 58.63 61.95 79.2 77.45 69.25 67.48 67.98 78.85 80.15 70.75 79.4 80.15 80.29
M2F 65.79 63.29 77.31 63.64 70.38 79.93 68.01 79.62 67.68 79.37 80.65 69.73 80.01 79.53 80.34
Tumour Core Ours 75.83 71.2 75.29 75.71 80.66 83.6 79.23 74.83 79.51 79.52 83.92 82.78 86.65 81.22 86.72
SPA 39.85 41.39 70.43 41.72 45.99 73.07 45.25 72.87 45.25 72.59 73.52 47.56 73.01 73.55 73.65
M3AE 37 38.41 75.8 44.22 78.09 45.2 79.36 78.16 41.71 48.12 79.14 80.06 47.63 79.31 79.91
mmFormer 40.08 72.19 38.89 37.23 73.11 73.06 40.64 42.27 43.65 75.56 43.34 81.74 73.36 75.31 73.4
RFNET 38.69 69.22 30.89 33.56 71.4 70.9 38.53 41.91 40.9 69.51 71.61 43.37 71.17 74.2 73.79
M2F 37.99 37.79 71.74 39.28 43.37 74.66 45.42 73.48 43.5 73.48 73.56 45.93 73.15 74.03 75.26
Enhancing tumour Ours 67.45 54.83 70.86 47.63 69.38 52.91 70.1 59.45 67.44 63.91 70.79 57.78 69.42 59.76 70.64
SPA 85.77 72.69 71.95 80.4 87.82 87.97 88.27 75.57 88.27 81.8 88.3 88.78 89.06 82.87 89.27
M3AE 87.78 74.69 74.91 84.43 76.09 84.48 89.63 84.4 88.64 88.91 84.04 89.29 88.58 88.45 88.26
mmFormer 84.09 72.85 73.37 85.6 85.97 76.93 87.09 86.09 87.55 87.94 88.36 88.16 88.74 85.96 89.03
RFNET 80.52 67.06 68.42 82.96 82.57 71.97 85.82 83.25 86 84.94 86.06 86.53 86.34 83.61 86.82
M2F 85.72 72.48 71.78 82.53 87.73 87.66 84.35 76.03 87.69 84.27 88.17 88.22 88.47 84.32 88.72
Whole tumour Ours 89.23 81.73 82.26 87.45 89.74 89.03 88.00 81.92 89.72 86.27 89.12 90.5 91.25 87.1 91.19
Figure 6: Generated images of our proposed method: slices across ventricles.
Figure 7: Generated images of our proposed method: slices across cerebral sulcus.
Figure 8: Generated images of our proposed method: slices across cerebellar hemisphere. Our method can generate defined cerebellar folia.

### v4

@@
-Personalize to generalize: Towards a universal medical multi-modality generalization through personalization
+Towards a Universal 3D Medical Multi-modality Generalization via Learning Personalized Invariant Representation
 Abstract
-Personalized medicine is a groundbreaking healthcare framework for the $21^{st}$ century, tailoring medical treatments to individuals based on unique clinical characteristics, including diverse medical imaging modalities.
-Given the significant differences among these modalities due to distinct underlying imaging principles, generalization in multi-modal medical image tasks becomes substantially challenging.
-Previous methods addressing multi-modal generalization rarely consider personalization, primarily focusing on common anatomical information.
-This paper aims to bridge multi-modal generalization with the concept of personalized medicine.
-Specifically, we propose a novel approach to derive a tractable form of the underlying personalized invariant representation $\mathbb{X}_{h}$ by leveraging individual-level constraints and a learnable biological prior.
-We demonstrate the feasibility and benefits of learning a personalized $\mathbb{X}_{h}$, showing that this representation is highly generalizable and transferable across various multi-modal medical tasks.
-Our method is rigorously validated on medical imaging modalities emphasizing both physical structure and functional information, encompassing a range of tasks that require generalization.
-Extensive experimental results consistently show that our approach significantly improves performance across diverse scenarios, confirming its effectiveness.
+Variations in medical imaging modalities and individual anatomical differences pose challenges to cross-modality generalization in multi-modal tasks.
+Existing methods often concentrate exclusively on common anatomical patterns, thereby neglecting individual differences and consequently limiting their generalization performance.
+This paper emphasizes the critical role of learning individual-level invariance, i.e., personalized representation $\mathbb{X}_{h}$, to enhance multi-modality generalization under both homogeneous and heterogeneous settings.
+It reveals that mappings from individual biological profile to different medical modalities remain static across the population, which is implied in the personalization process.
+We propose a two-stage approach: pre-training with invariant representation $\mathbb{X}_{h}$ for personalization, then fine-tuning for diverse downstream tasks.
+We provide both theoretical and empirical evidence demonstrating the feasibility and advantages of personalization, showing that our approach yields greater generalizability and transferability across diverse multi-modal medical tasks compared to methods lacking personalization.
+Extensive experiments further validate that our approach significantly enhances performance in various generalization scenarios.
 1 Introduction
-Figure 1: A diagram of medical modalities and individual differences. Individual variations may be significant and warrant further research attention from the medical intelligence community.
-Personalized medicine represents a transformative framework for $21^{st}$ century healthcare, tailoring medical treatments to each patient’s unique characteristics [45, 24, 4].
-This approach necessitates diverse information, including clinical data such as radiological images.
-Three-dimensional medical images, generated through specialized techniques and radiopharmaceuticals, excel at highlighting specific anatomical features.
-Collectively, different medical image modalities provide a comprehensive view of a patient’s structural and functional characteristics.
-However, this distinction between medical modalities creates significant generalization challenges in medical image analysis, especially when certain modalities may be inaccessible due to an individual’s financial constraints or physical limitations, thereby complicating the effectiveness of personalized medicine.
+Figure 1: A diagram of medical modalities and individual differences. These variations are significant and deserve further study by the medical intelligence community.
+Three-dimensional radiological medical images, generated through specialized techniques and radiopharmaceuticals, excel at highlighting specific physiological features, collectively providing a comprehensive view of a patient’s structural and functional characteristics.
 As illustrated in Fig.
-1, contemporary research in medical intelligence is mainly concentrated on structural modalities that depict physical anatomy.
-This includes Magnetic Resonance Imaging (MRI) scans [53], which use strong magnetic fields and radiofrequency currents yielding distinct sequences, and Computed Tomography (CT) scans [33, 51], which employ X-rays to measure its attenuation.
-Other studies [49] focus on the functional modalities associated with biochemistry, such as Positron Emission Tomography (PET) scans.
-PET scans are expensive functional imaging scans that employ radiotracers emitting gamma rays to visualize and measure metabolic processes.
-The differing imaging principles result in substantial modality gaps, presenting a critical challenge for model generalization.
-For clarity, we categorize these modalities for generalization tasks into two types: homogeneous generalization, which pertains to generalizing within structural or functional modalities (e.g., T1, T2, T1ce, and Flair in MRI, as shown in Fig.
-1); and heterogeneous generalization, which involves generalizing across both structural and functional modalities, such as CT and PET.
-As each person is fundamentally different from the average of the population [45], the concept of multi-modality generalization needs comprehensive discussion within the scope of personalization, an area scarcely addressed in previous research.
-An ideally well-generalized personalized medical model across modalities should (1) provide additional insights derived from all modalities when only a subset is accessible and (2) seamlessly transfer across domains while maintaining the capacity to achieve (1).
-While pre-training can significantly enhance downstream generalization, some recent approaches concentrate on learning common physical anatomy invariance at the class level [23], which may overlook the individual variations.
-Another line of research [40, 46, 23] primarily addresses the transferability of single-modal tasks and may not be suitable for multi-modal scenarios.
-In addition, most research on homogeneous generalization for medical tasks focuses on structural sequences of MRI or CT, employing strategies such as cross-modality transfer [30, 25, 51] or targeting challenges like missing modality segmentation [31, 6, 36, 37, 51].
-Heterogeneous generalization presents a greater challenge due to the disparities between structural and functional information.
-Despite its significance, very few efforts [34] have addressed heterogeneous generalization, mainly focusing on one-directional modality transfer (e.g., PET to CT or MRI), and rarely exploring the model’s generalizability and transferability for other tasks in this context.
-To address multi-modality generalization issue for personalization under both homogeneous and heterogeneous settings, we formally introduce the concept of the personalized invariant representation for multi-modal generalization, denoted as $\mathbb{X}_{h}$, and its constraints as outlined in Hypothesis 3.1.
-Furthermore, personalized invariant $\mathbb{X}_{h}$, which learns aggregated biological information from all possible modalities specific to the individual, is likely to enhance performance across various medical tasks for that person.
-Building on this hypothesis, this paper proposes a general approach aimed at enhancing the generalization of various medical imaging tasks through personalization.
-Specifically, our method constructs an approximation of $\mathbb{X}_{h}$ using the learnable biological prior knowledge $\mathbb{O}$, via decomposition, invariance, and equivariance constraints during pre-training (refer to Sec.
-3.2).
-The learned approximation of $\mathbb{X}_{h}$ can then be utilized to enhance performance in downstream generalization tasks, irrespective of whether a domain gap exists between the pre-training data and downstream data.
-Importantly, this paper demonstrates that obtaining a personalized invariant representation, $\mathbb{X}_{h}$, is feasible through our approach, and such invariance leads to generalization improvements across various medical tasks.
+1, current research in medical intelligence mainly targets structural modalities like Magnetic Resonance Imaging (MRI) [58] and Computed Tomography (CT) [36, 56] scans.
+Other studies [54] focus on the functional modalities associated with biochemistry, like Positron Emission Tomography (PET) scans.
+Notably, all modalities exhibit significant individual variations.
+Individuals differ fundamentally from population averages [50], making anatomical and metabolic variations evident.
+In clinical settings, financial or physical limitations may restrict access to certain modalities, highlighting the need for generalization in medical image analysis across different modalities and individuals.
+The generalization faces significant challenges due to modality and individual variations.
+We classify generalization tasks into two types: homogeneous generalization within structural or functional modalities, e.g., MRI sequences Fig.
+1; and heterogeneous generalization across both structural and functional modalities, such as CT and PET scans.
+A well-generalized medical model should integrate insights from all available modalities to support downstream tasks, relying on condition invariance [10, 43].
+Some methods enhance generalization through class-level anatomy invariance [25], which may not fit models based on functional biochemistry and may degrade transferability for unseen individual domains.
+Single-modal task transfer research [44, 51, 25] might not translate to multi-modal scenarios.
+Homogeneous generalization in medical tasks often employs cross-modality transfer [32, 27, 56] or tackles missing modalities [34, 7, 40, 41, 56] in MRI or CT, while heterogeneous generalization is insufficiently explored.
+Furthermore, these limited efforts often focus on modality or class-level invariance, neglecting individual differences and impeding a model’s capacity to generalize effectively across patient populations and modalities.
+Figure 2: A diagram of patient domains and medical modalities linked to our method.
+In the Pre-training stage, static mappings between domains and modalities are implicitly learned.
+The fine-tuning stage addresses gaps between seen and unseen patients.
+This paper reveals that both homogeneous and heterogeneous settings for multi-modality generalization can be tackled within the framework of personalization, i.e., learning the personal-level invariant representation.
+Based on medical imaging principles, mappings from an individual’s identical biological profile to various modalities remain static across individuals.
+Learning personal-level invariant representations, which apply across multiple modalities, enhances cross-modality generalization and transferability to novel individuals.
+To formalize this approach, the hypothesis of a personalized invariant representation is introduced, denoted as $\mathbb{X}_{h}$, which exists for each individual in multi-modal generalization.
+Hypothesis  3.1 rigorously details the constraints and properties of $\mathbb{X}_{h}$.
+Building on this hypothesis, a general approach termed PUIR is proposed to augment the generalization of various medical imaging tasks through personalization.
+As shown in Fig.
+2, our approach consists of two stages: (1) The pre-training stage constructs an approximation of $\mathbb{X}_{h}$ using the learnable biological prior knowledge $\mathbb{O}$, via decomposition, invariance, and equivariance constraints during pre-training (refer to Sec.
+3.1.1), focusing on individual-level modality generalization.
+(2) The fine-tuning stage, employs standard downstream training methods to adapt the modality encoders derived from the first stage to mitigate the gap between seen and unseen individuals.
+A key distinction between our method and previous pre-training strategies lies in our specific focus on enhancing generalization across various medical modalities and individuals.
+Importantly, this paper theoretically demonstrates that obtaining a personalized invariant representation, $\mathbb{X}_{h}$, is feasible through our approach, and such invariance leads to generalization improvements across various medical tasks (see Sec.
+3.3).
 To validate our methodology, we conduct experiments on modality transfer and missing modality segmentation tasks, addressing not only the homogeneous generalization of MRI but also the rarely explored heterogeneous generalization, such as generalization between PET and CT.
-Our findings reveal that our approach successfully captures comprehensive, personalized information even when only partial modalities are available for a given individual (see Fig.
-3).
+Our findings reveal that our approach successfully captures comprehensive, personalized information even when only partial modalities are available for a given individual.
 Moreover, extensive experiments on both homogeneous (Sec.
-4) and heterogeneous (Secs.
-5 and 6) generalization demonstrate that our approach can be adapted for downstream tasks and surpasses current state-of-the-art (SOTA) methods in multiple tasks, validating its superiority.
-We will publicly release our code, checkpoints, and data splits upon acceptance.
+4.1) and heterogeneous (Secs.
+4.2 and 4.3) generalization demonstrate that our approach can be adapted for downstream tasks and surpasses current state-of-the-art (SOTA) methods in multiple tasks, validating its superiority.
+Our code and data splits are available at https://github.com/zhaorui-tan/PUIR_ICCV25.
 2 Related work
 Medical generalization tasks currently concentrates on homogeneous generalization, introducing tasks, such as modality transfer and missing modality segmentation for structural modalities — Flair, T1, T2, and T1ce of MRI- in brain tumor segmentation [58], or between MRI and CT [56] for modality transfer.
-[34] propose an approach for heterogeneous generalization in terms of modality transfer, but only tailored for transferring PET to CT.
+[38] proposes an approach for heterogeneous generalization in terms of modality transfer, but only tailored for transferring PET to CT.
 In terms of tasks, most current studies focus on one specific generalization task, either segmentation [6, 47, 14, 57, 48, 33] or modality transfer [24, 60, 15, 39, 28, 32, 42, 13, 36, 27, 52, 56].
 This paper aims to develop an approach that is feasible for different downstream tasks under both homogeneous and heterogeneous modality generalization.
-Our approach aims to learn the $\mathbb{X}_{h}$ through pre-training; we list related medical pre-training work [40, 46, 8, 23] here.
+Our approach aims to learn the $\mathbb{X}_{h}$ through pre-training; we list related medical pre-training work [44, 51, 9, 25, 37] here.
 A notable work among them is [25], which extracts class-specific anatomical invariance.
@@
 The generalization in medical images also connects to alignment in multi-domain generalization for natural images [16, 30, 29, 22, 43].
-Please refer to Appendix A for a detailed literature review.
+Please refer to Appendix B for a detailed literature review.
+Figure 3: Left: Overall framework of learning $\mathbb{X}_{h}$ during the pre-training stage.
+Right: Diagrams of differences between previous methods of learning $\mathbb{Z}_{h}$ and our proposed method of learning $\mathbb{X}_{h}$.
 3 Learning personalized invariant representation for medical generalization
-Preliminaries.
-In this paper, we denote the encoder as $\mathcal{E}$ and its corresponding decoder as $\mathcal{D}$.
-For an individual human being $h\in\mathcal{H}$, the corresponding medical images are represented as $X_{h}={X^{i}_{h},X^{j}_{h},\dots,X^{k}_{h}}$, where $i,j,\dots,k\in\mathcal{M}$, and $\mathcal{M}$ represents the set of all possible modality combinations.
-We denote the intermediate features produced by $\mathcal{E}(X_{h})$ and $\mathcal{E}(X^{i}_{h})$ as $x_{h}$ and $x^{i}_{h}$, respectively.
-The final layer features from the encoder are represented as $z_{h}$ and $z^{i}_{h}$.
-The learned approximation of $\mathbb{X}_{h}$ is denoted as ${\mathbb{X}_{h}}^{\prime}$.
-Additionally, we define the geometric warping function $\phi^{i}\in\Phi$, where $\phi^{i}(X^{i}_{h})\in\mathcal{X}$, and $\Phi$ denotes the set of all possible geometric warping functions.
-Finally, $I(\cdot;\cdot)$ and $P(\cdot)$ represent mutual information and probability distribution, respectively.
-Before addressing the problem for both homogeneous and heterogeneous generalization, we introduce the $\mathbb{X}_{h}$ Hypothesis for medical imaging:
+Notations.
+The encoder is denoted to $\mathcal{E}$, while its associated decoder is defined as $\mathcal{D}$.
+An individual human entity, denoted as $h\in\mathcal{H}$, with corresponding medical images $X_{h}={X^{i}_{h},X^{j}_{h},\dots,X^{k}_{h}}$, whereby $i,j,\dots,k\in\mathcal{M}$, and $\mathcal{M}$ cover all conceivable modality combinations.
+Intermediate features generated by $\mathcal{E}(X_{h})$ and $\mathcal{E}(X^{i}_{h})$ are designated as $x_{h}$ and $x^{i}_{h}$, respectively.
+The encoder’s final layer features are $z_{h}$ and $z^{i}_{h}$.
+The learned approximation of $\mathbb{X}_{h}$ is expressed as ${\mathbb{X}_{h}}^{\prime}$.
+The geometric warping function is $\phi^{i}\in\Phi$, with $\phi^{i}(X^{i}_{h})\in\mathcal{X}$ and $\Phi$ as the ensemble of all possible warping functions.
+Conclusively, $I(\cdot;\cdot)$ and $P(\cdot)$ signify mutual information and probability distribution, respectively.
+Personalized invariant representation hypothesis.
+Before addressing the problem for both homogeneous and heterogeneous generalization, we introduce the personalized invariant representation hypothesis, termed as $\mathbb{X}_{h}$ Hypothesis, for medical imaging:
 Hypothesis 3.1 ($\mathbb{X}_{h}$ Hypothesis).
-Consider the set $\mathcal{M}$ of all possible modality combinations and the set $\Phi$ of all possible geometric transformations (e.g., SO(3)) transformations, for example, rotations corresponding to different poses of the person.
-There exists a personalized invariant representation $\mathbb{X}_{h}$ for an individual from the population $h\in\mathcal{H}$, which can be decomposed into modality-specific images $X^{i}_{h}$ given a modality combination $i\in\mathcal{M}$ as a condition:
+Consider the set $\mathcal{M}$ of all possible modality combinations and the set $\Phi$ of possible geometric transformations (e.g., SO(3)), such as rotations corresponding to different postures of an individual.
+For a given individual from the population $h\in\mathcal{H}$, there exists a personalized invariant representation $\mathbb{X}_{h}$, which can be decomposed into modality-specific images $X^{i}_{h}$ conditioned on a modality combination $i\in\mathcal{M}$.
+This relationship is formally expressed as:
 $X^{i}_{h}=\mathbb{X}_{h}|i;\;i\in\mathcal{M},h\in\mathcal{H},\quad\textit{s.t.,}\mathbb{X}_{h}\perp\!\!\!\perp\mathcal{M},\Phi.$ (1)
@@
 The overall learning framework for $\mathbb{X}_{h}$ is illustrated on the left-hand side of Fig.
-1.
+3.
 Data from each modality are encoded by $\mathcal{E}$, and the encoded features are used to retrieve knowledge from the learnable biological prior $\mathbb{O}$.
@@
 By applying constraints of decomposition, equivariance, and invariance on the fused features, we approximate $\mathbb{X}_{h}$ effectively.
-Figure 2: Left: Overall framework of learning $\mathbb{X}_{h}$.
-Right: Diagrams of differences between previous methods of learning $\mathbb{Z}_{h}$ and our proposed method of learning $\mathbb{X}_{h}$.
 As illustrated in Fig.
@@
 The reason is that the transferred targets become constrained by the learned modal-dependent features, potentially limiting their broader applicability.
-3.1 Using global prior for better personalized invariant representation
-To learn a better approximation of $\mathbb{X}_{h}$, we leverage a global biological prior, denoted as $\mathbb{O}$.
+3.1 Pre-training stage
+To learn a better approximation of $\mathbb{X}_{h}$, we leverage a learnable biological prior, denoted as $\mathbb{O}$.
 If $\mathbb{O}$ can be learned, representations from any modality can complete themselves by retrieving the missing knowledge from $\mathbb{O}$, forming a better approximation of $\mathbb{X}_{h}$.
@@
 As shown in Fig.
-1, the representation $z^{i}_{h}$ retrieves its missing knowledge from $\mathbb{O}$ via attention: ${z^{i}_{h}}^{\prime}:=attn(query:z^{i}_{h},key:\mathbb{O},value:\mathbb{O})$.
+3, the representation $z^{i}_{h}$ retrieves its missing knowledge from $\mathbb{O}$ via attention: ${z^{i}_{h}}^{\prime}:=attn(query:z^{i}_{h},key:\mathbb{O},value:\mathbb{O})$.
 The original representation and the retrieved knowledge are then fused through convolution: $\mathbb{X}^{i}_{h}:=conv({z^{i}_{h}}^{\prime},z^{i}_{h})$.
@@
 The details of these constraints are discussed in Sec.
-3.2.
-3.2 Constraints of equivariance, invariance, and decomposition
+3.1.1.
+3.1.1 Constraints of equivariance, invariance, and decomposition
 Contrastive learning.
@@
 Therefre, if $\phi^{i}$ can be extracted from the last-layer output $z^{i}_{h}$, it can also be extracted from the $\phi^{i}(x^{i}_{h})$ from the previous layers:
-$\min_{\mathcal{D},\mathcal{F}}Dis(\phi^{i},\mathcal{F}(z^{i}_{h})),$ (4)
+$\min\nolimits_{\mathcal{D},\mathcal{F}}Dis(\phi^{i},\mathcal{F}(z^{i}_{h})),$ (4)
 where $\mathcal{F}:\mathcal{F}(z^{i}_{h})\to{\phi^{i}}^{\prime}$ extracts the geometric transformation and $Dis(\dot{,}\dot{)}$ denotes the distance measurement between ${\phi^{i}}^{\prime}$ and $\phi^{i}$.
@@
 Under this case, $\Phi_{R}$ is defined as rotations at $[0,90,180,270]$ degrees along the z-axis, and $\phi_{r}^{i}\in\Phi_{R}$ is the ground truth rotation categories.
-$\mathcal{F}{z^{i}_{h}}$ produces the softmax probabilities of rotation categories.
-The loss is in the form of:
+$\mathcal{F}{z^{i}_{h}}$ produces the softmax probabilities of rotation categories, predicting which kind of rotation is applied, and loss is in the form of:
 $\mathcal{L}_{equ}=-\sum\nolimits_{r=1}^{|\Phi_{R}|}\phi_{r}^{i}\log\mathcal{F}(z^{i}_{h}).$ (5)
-Task T1→T2 T2 → Flair T1 → T1ce
-Method PSNR↑ NMSE↓ SSIM↑ PSNR↑ NMSE↓ SSIM↑ PSNR↑ SSIM↑
-2D
-Pix2Pix [22] 24.624 0.109 0.874 24.82 0.0250 0.846 27.05 0.858
-CycleGAN [54] 23.535 0.155 0.837 23.418 0.164 0.825 30.13 0.906
-NICEGAN [7] 23.721 0.148 0.840 23.643 0.148 0.829 - -
-GcGAN [13] - - - 29.98 - 0.917 25.98 0.872
-CUT [35] - - - 23.54 - 0.819 26.27 0.846
-RegGAN [26] 24.884 0.094 0.881 24.576 0.112 0.852 31.36 0.930
-ResViT [10] 25.578 0.088 0.895 24.825 0.108 0.861 31.46 0.932
-Diffusion [11] - - - 31.98 - 0.930 29.22 0.921
-MD-Diff [47] - - - 30.76 - 0.934 33.08 0.948
-3D
-Pix2Pix 23.740 0.138 0.835 23.508 0.152 0.822 - -
-CycleGAN 25.181 0.097 0.887 24.602 0.113 0.854 - -
-EaGAN [50] 24.884 0.094 0.881 24.576 0.112 0.852 - -
-MS-SPADE [25] 25.818 0.079 0.904 25.074 0.098 0.867 26.119 0.912
-Ours 30.756 0.065 0.944 32.224 0.046 0.941 34.547 0.955
-Table 1: Modality transfer results of MRI on BRATS23: Comparison between previous methods and our method.
-The best results are highlighted in blue.
-Results denoted with are gained from [25], while the results denoted with are gathered from [47].
-Target T1 T1ce T2 Flair
-Source PSNR↑ NMSE↓ SSIM↑ PSNR↑ NMSE↓ SSIM↑ PSNR↑ NMSE↓ SSIM↑ PSNR↑ NMSE↓ SSIM↑
-SwinUNETR 32.815 0.092 0.941 31.655 0.202 0.912 24.650 0.361 0.857 27.593 0.202 0.883
-MS-SPADE 29.001 0.055 0.942 26.119 0.078 0.912 25.818 0.103 0.904 24.842 0.113 0.859
-T1 Ours 43.472 0.003 0.996 34.547 0.045 0.955 30.756 0.065 0.944 31.693 0.049 0.937
-SwinUNETR 32.456 0.100 0.929 33.001 0.156 0.926 25.125 0.366 0.859 27.699 0.211 0.882
-MS-SPADE 26.228 0.076 0.922 28.759 0.060 0.937 25.990 0.092 0.907 25.204 0.092 0.881
-T1ce Ours 34.077 0.020 0.962 46.663 0.003 0.996 30.775 0.063 0.942 32.224 0.046 0.941
-SwinUNETR 30.102 0.171 0.896 30.354 0.283 0.883 26.831 0.268 0.887 27.234 0.242 0.872
-MS-SPADE 25.422 0.085 0.908 25.234 0.087 0.895 29.230 0.048 0.942 25.074 0.098 0.867
-T2 Ours 32.646 0.028 0.955 33.857 0.051 0.949 43.653 0.006 0.991 32.224 0.046 0.941
-SwinUNETR 31.371 0.135 0.916 31.285 0.240 0.905 25.579 0.338 0.867 29.092 0.148 0.923
-MS-SPADE 25.186 0.090 0.905 25.899 0.094 0.906 26.146 0.086 0.913 28.608 0.058 0.938
-Flair Ours 32.752 0.026 0.951 33.471 0.055 0.944 30.571 0.068 0.940 43.624 0.004 0.995
-Table 2: Modality transfer results of MRI on BRATS23: Comparison between the previous method and ours for transfer between all four modalities.
-The averaged results of metrics for all validation samples are listed.
-The best results are highlighted in blue.
-Please refer to Appendix 9 for the standard derivations according to each result.
 Decomposition constraint.
@@
 An intuitive approach is reconstructing all possible modalities by using $\mathbb{X}_{h}$, whose objective can be formed as:
-$\mathcal{L}_{decom}\!=\!\sum_{1}^{|\mathcal{M}|}\left\|{\phi^{i}}^{-1}\!\left(\mathcal{D}\left({\mathbb{X}_{h}}^{\prime}|\phi^{i}(x^{i}_{h})\right)\right),X_{h}\right\|^{2},$ (6)
+$\mathcal{L}_{decom}\!=\!\sum\nolimits_{1}^{|\mathcal{M}|}\left\|{\phi^{i}}^{-1}\!\left(\mathcal{D}\left({\mathbb{X}_{h}}^{\prime}|\phi^{i}(x^{i}_{h})\right)\right),X_{h}\right\|^{2},$ (6)
 where $i\in\mathcal{M}$ and $\phi^{i}(x^{i}_{h})$ represents intermediate representations produced during $\mathcal{E}(\phi^{i}(X^{i}_{h}))$ and $X_{h}$ denotes all possible modalities.
@@
 where the weight of each loss is omitted here.
+Task T1→T2
+Method PSNR↑ NMSE↓ SSIM↑
+2D
+Pix2Pix [24] 24.624 0.109 0.874
+CycleGAN [60] 23.535 0.155 0.837
+NICEGAN [8] 23.721 0.148 0.840
+RegGAN [28] 24.884 0.094 0.881
+ResViT [12] 25.578 0.088 0.895
+3D
+Pix2Pix 23.740 0.138 0.835
+CycleGAN 25.181 0.097 0.887
+EaGAN [55] 24.884 0.094 0.881
+Ours (PUIR) 30.756 0.065 0.944
+Task T2 → Flair
+Method PSNR↑ NMSE↓ SSIM↑
+2D
+Pix2Pix [24] 24.82 0.0250 0.846
+CycleGAN [60] 23.418 0.164 0.825
+NICEGAN [8] 23.643 0.148 0.829
+RegGAN [28] 24.576 0.112 0.852
+ResViT [12] 24.825 0.108 0.861
+Diffusion [13] 31.98 - 0.930
+MD-Diff [52] 30.76 - 0.934
+3D
+Pix2Pix 23.508 0.152 0.822
+CycleGAN 24.602 0.113 0.854
+EaGAN [55] 24.576 0.112 0.852
+Ours (PUIR) 32.224 0.046 0.941
+Task T1 → T1ce
+Method PSNR↑ SSIM↑
+2D
+Pix2Pix [24] 27.05 0.858
+CycleGAN [60] 30.13 0.906
+GcGAN [15] 25.98 0.872
+CUT [39] 26.27 0.846
+RegGAN [28] 31.36 0.930
+ResViT [12] 31.46 0.932
+Diffusion [13] 29.22 0.921
+MD-Diff [52] 33.08 0.948
+3D
+Ours (PUIR) 34.547 0.955
+Table 1: Modality transfer results of MRI on BRATS23: Comparison between previous one-to-one modality transfer methods and our method.
+The best results are highlighted in blue.
+Results denoted with are gained from [27]; with are gathered from [52].
 3.1.2 The connection between the constraints and global biological prior
@@
 This implicitly constrains $\mathbb{O}$ to contain comprehensive biological information, including other potential geometric transformations, thereby improving the $\mathbb{X}_{h}^{\prime}$ through $\mathbb{O}$ in approximating $\mathbb{X}_{h}$ and enhancing the robustness of ${\mathbb{X}_{h}}^{\prime}$.
-3.3 Application of different modalities and tasks
+3.2 Fine-tuning stage
 After pre-training with the loss function $\mathcal{L}_{pre}$, the model is then utilized for downstream tasks such as segmentation or generation.
@@
 Further empirical details on how $\mathbb{X}_{h}$ is leveraged for homogeneous and heterogeneous generalization are provided in Secs.
-4 and 5.
-4 Homogeneous generalization: MRI
-Figure 3: Visualization of the efficacy of prior $\mathbb{O}$.
-Displayed are the generated modalities on the input Flair modality of a testing sample on the BTATS21 dataset.
-Columns show: the generated images of the model (1) without prior $\mathbb{O}$ and (2) with prior $\mathbb{O}$ are aligned with (3) the GT images.
-Typically, the differences between without and with prior $\mathbb{O}$ (the (2)-(1) column) are visualized to compare with the differences between without $\mathbb{O}$ and GT (the (3)-(1) column).
-Red and blue refer to the positive (accomplishment) and negative (refinement) values of the differences, respectively.
-This section demonstrates that our approach enhances the homogeneous generalization across structural modalities in MRI.
-To validate that our method captures personalized information, especially anatomical structure features, during the pre-training stage, we first apply it to modality transfer tasks.
-Next, we adapt the pre-trained model for the downstream missing modality segmentation task.
-Experimental results indicate that our approach outperforms state-of-the-art (SOTA) methods in both tasks, thereby supporting the $\mathbb{X}_{h}$ Hypothesis and confirming the effectiveness of our method.
-4.1 Pre-training for modality transfer
-Modality transfer tasks focus on converting medical images from multiple modalities to other modalities.
-We test our approach on the structural modalities of MRI.
-This task aligns seamlessly with our pre-training objective, and the quality of the generated modalities serves as a validation of the anatomical knowledge captured by our pre-trained model.
-Importantly, our approach can generate all modalities without knowing the exact modalities where the input form, as the learned representation ${\mathbb{X}_{h}}^{\prime}$ encompasses comprehensive information across all possible modalities.
+4.1 and 4.2.
+3.3 Theoretical analysis
+Consider a mapping $g:\mathcal{H}\to\mathcal{M}$ that extracts modalities from the individual biological profile, where $g\in\mathcal{G}$ represents a set of candidate hypotheses.
+In modality generalization, our objective is to guarantee that the model effectively generalizes across diverse modalities $\mathcal{M}$ for previously unseen patients.
+To achieve this, the learning bound must account for the relationships between seen $\mathcal{H}^{\mathcal{S}}$ and unseen patients $\mathcal{H}^{\mathcal{U}}$.
+Inspired by the theoretical framework of domain adaptation (DA) risk bounds in [4], we derive the following bound for medical cross-modality generalization:
+Theorem 3.2.
+Let $R^{\mathcal{S}}$ and $R^{\mathcal{U}}$ denote the generalization errors on the modalities from the seen patient domain $\mathcal{D}^{\mathcal{S}}$ and the unseen patient domain $\mathcal{D}^{\mathcal{U}}$, respectively.
+For a given hypothesis $g\in\mathcal{G}$, the overall unseen risk is bounded by:
+$\begin{split}R^{\mathcal{U}}(g)\leq R^{\mathcal{S}}(g)+d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{D}^{\mathcal{S}},\mathcal{D}^{\mathcal{U}})+\lambda,\end{split}$ (9)
+where $d_{\mathcal{G}\Delta\mathcal{G}}$ is the $\mathcal{G}\Delta\mathcal{G}$-divergence between $\mathcal{D}^{\mathcal{S}}$ and $\mathcal{D}^{\mathcal{U}}$, and joint hypothesis $\lambda$ quantifies the inherent difficulty of aligning the seen and unseen domains (can be omitted).
+The bound emphasizes the importance of minimizing the divergence between domains and achieving a low risk on the seen domain to ensure effective generalization to unseen domains.
+For a well-trained model on $R^{\mathcal{S}}(g)$ is minimized.
+the key focus shifts to $d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{D}^{\mathcal{S}},\mathcal{D}^{\mathcal{U}})$.
+Considering modalities, $d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{D}^{\mathcal{S}},\mathcal{D}^{\mathcal{U}})$ is defined as follows:
+$\begin{split}d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{D}^{\mathcal{S}},\mathcal{D}^{\mathcal{U}})\!:=d_{\mathcal{G}\Delta\mathcal{G}}([\mathcal{H}^{\mathcal{S}}\oplus\mathcal{M}|\mathcal{H}^{\mathcal{S}}],[\mathcal{H}^{\mathcal{U}}\oplus\mathcal{M}|\mathcal{H}^{\mathcal{U}}]),\end{split}$
+where $\oplus$ denotes that each domain combines the patient domain with its derived modalities. Upon rearranging, it follows that:
+$d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{H}^{\mathcal{S}},\mathcal{H}^{\mathcal{U}})+d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{M}|\mathcal{H}^{\mathcal{S}},\mathcal{M}|\mathcal{H}^{\mathcal{U}}).$ (10)
+Our approach minimizes $d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{M}|\mathcal{H}^{\mathcal{S}},\mathcal{M}|\mathcal{H}^{\mathcal{U}})$ by initially learning $\mathbb{X}_{h}$ for each patient at the pre-training stage and then employing a fine-tuning stage to minimize $d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{H}^{\mathcal{S}},\mathcal{H}^{\mathcal{U}})$.
+As shown in Fig.
+2, while each modality uniquely visualizes a patient’s biological profile, the underlying generation principles remain unchanged.
+Thus, the mapping $g_{m}:\mathcal{H}\to\mathcal{X}_{m}$ that derives the modality-specific visualizations $m\in\mathcal{M}$ from the individual biological profile is consistent across individuals.
+Since direct access to a person’s anatomical structure is infeasible, we propose learning an invariant representation $\mathbb{X}_{h}$ for $h\in\mathcal{H}$ with a biological prior, allowing decomposition into various modalities.
+To ensure $\mathbb{X}_{h}$ effectively captures anatomical features, we enforce equivariance constraints to preserve geometric information.
+This learned representation augments the quality of $g_{m}$, improving its generalization ability by reducing $d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{M}|\mathcal{H}^{\mathcal{S}},\mathcal{M}|\mathcal{H}^{\mathcal{U}})$.
+In the second fine-tuning stage, it is only necessary to mitigate $d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{H}^{\mathcal{S}},\mathcal{H}^{\mathcal{U}})$, thus facilitating the transfer process.
+Previous approaches ignore personal variations, opting instead to focus on minimizing the gap between modalities, as noted in $d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{M}^{\mathcal{S}},\mathcal{M}^{\mathcal{U}})$.
+To facilitate a more effective comparison with our approach, we simplify this issue by positing that $d_{\mathcal{G}\Delta\mathcal{G}}$ is quantifiable via Kullback-Leibler (KL) divergence.
+Consequently, previous method aimed at minimizing $d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{M}^{\mathcal{S}},\mathcal{M}^{\mathcal{U}})$ can be converted as follows:
+$\begin{split}&d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{M}^{\mathcal{S}},\mathcal{M}^{\mathcal{U}}):=KL(\mathcal{M}^{\mathcal{S}}\|\mathcal{M}^{\mathcal{U}})\\ =&KL(\mathcal{H}^{\mathcal{S}}\|\mathcal{H}^{\mathcal{U}})+E_{\mathcal{H}^{\mathcal{S}},\mathcal{H}^{\mathcal{U}}}[KL(\mathcal{M}^{\mathcal{S}}|\mathcal{H}^{\mathcal{S}}\|\mathcal{M}^{\mathcal{U}}|\mathcal{H}^{\mathcal{U}})]\\ \geq&KL(\mathcal{M}^{\mathcal{S}}|\mathcal{H}^{\mathcal{S}}\|\mathcal{M}^{\mathcal{U}}|\mathcal{H}^{\mathcal{U}})=:d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{H}^{\mathcal{S}},\mathcal{H}^{\mathcal{U}})\;(\textbf{Ours}).\end{split}$
+Our methodology narrows the gap identified in Eq. 10 compared to earlier methods, thereby demonstrating enhanced generalization and transferability capabilities.
+4 Experiments
+Overall experiment arrangement.
+Our experiments examine various generalization scenarios in two stages: generation pre-training and downstream fine-tuning.
+We assess pre-training by comparing generation quality to SOTA medical generation methods and evaluate fine-tuning by comparing with SOTA methods for each task, including benchmarking against SOTA medical self-supervised pre-training methods.
+More results and analysis are in Appendix G.
+4.1 Homogeneous generalization: MRI
+From scratch PSNR↑ NMSE↓ SSIM↑
+SwinUNETR 29.1776 0.2196 0.8961
+MS-SPADE 26.4225 0.0822 0.9086
+Ours (PUIR) ($96\times 96\times 96$) 35.4378 0.0362 0.9587
+Ours (PUIR) ($192\times 192\times 192$) 33.7418 0.0481 0.9502
+Table 2: Modality transfer results of MRI on BRATS23: Comparison between the previous one-to-all modality transfer methods and ours for transfer between all four modalities.
+The averaged results of metrics for all validation samples across all modalities are listed.
+The best results are highlighted in blue.
+Please refer to Appendix Tab.
+10 for more detailed results.
+4.1.1 Pre-training stage: Modality transfer
+Experimental settings.
+Following previous methods [27], we utilize the multi-modal brain tumor segmentation challenge 2023 (BRATS23) dataset [1, 35, 3, 2].
+BRATS23 includes four structural MRI modalities (T1, T1ce, T2, and FLAIR) for each individual.
+Our model is tested on the BRATS23 validation set, which contains these four modalities for 219 individuals.
+We evaluate the quality of synthesis using peak signal-to-noise ratio (PSNR), normalized mean squared error (NMSE), and structural similarity index (SSIM) [53].
+To provide comprehensive results, we separately compare the translation results for T1 $\to$ T2 and T2 $\to$ FLAIR, as some previous methods are only capable of single-modality transfer.
+These include both 2D and 3D generation methods, as shown in Tab.
+1.
+Additionally, we employed SwinUNTER for multi-modality translation comparisons.
+All evaluations were performed on 3D volumes; for the 2D methods, synthesized target images were stacked to form a 3D volume for comparison.
+Please refer to model training details to Appendix F.
 All settings FM MN=1 MN=2 MN=3
@@
 M2F [42] 73.69 6.83 80.34 77.48 5.19 74.17 6.88 67.51 6.63
-Ours 79.78 4.55 86.72 83.64 2.29 79.56 3.28 74.51 3.87
+SwinUNETR pre-training 79.19 4.24 84.69 82.59 3.43 79.19 2.69 74.41 3.54
+Ours (PUIR) 79.78 4.55 86.72 83.64 2.29 79.56 3.28 74.51 3.87
 Enhancing tumor
@@
 M2F 58.84 16.58 75.26 66.67 13.83 58.99 16.19 46.70 15.97
-Ours 63.49 7.58 70.64 64.44 6.62 63.87 11.45 60.19 11.31
+SwinUNETR pre-training 59.47 6.25 61.38 60.61 3.43 60.85 6.10 55.79 9.14
+Ours (PUIR) 63.49 7.58 70.64 64.44 6.62 63.87 11.45 60.19 11.31
 Whole tumor
@@
 M2F 83.88 5.79 88.72 87.30 1.99 84.62 2.57 78.13 7.81
-Ours 87.63 3.25 91.19 89.49 1.82 87.45 1.03 85.17 3.93
-Table 3: Missing modality segmentation results of MRI on BRATS18: Num denotes the number of missing modalities for different settings.
-We report mean and standard deviations of DICE results of all experimental combinations under the same Num.
-The best results are highlighted in blue.
+SwinUNETR pre-training 86.55 3.49 89.22 88.18 1.38 86.69 3.15 84.04 4.91
+Ours (PUIR) 87.63 3.25 91.19 89.49 1.82 87.45 1.03 85.17 3.93
+Table 3: Missing modality segmentation results of MRI on BRATS18: Comparisons between previous SOTA methods and ours.
+MN: Number of how many missing modalities; FM: Full Modality.
+We report mean and standard deviations of DICE results under the MN.
+The best DICE results are highlighted in blue.
 Please refer to Appendix Tab.
-10 for detailed results of each setting.
-MN: Missing Num; FM: Full Modality.
+11 for more details.
 SSIM↑ PSNR↑
@@
 The best and second results are highlighted in blue and cyan, respectively.
-Experimental settings.
-Following previous methods [25], we utilize the multi-modal brain tumor segmentation challenge 2023 (BRATS23) dataset [1, 32, 3, 2].
-BRATS23 includes four structural MRI modalities (T1, T1ce, T2, and FLAIR) for each individual.
-Our model is tested on the BRATS23 validation set, which contains these four modalities for 219 individuals.
-We evaluate the quality of synthesis using peak signal-to-noise ratio (PSNR), normalized mean squared error (NMSE), and structural similarity index (SSIM) [48].
-To provide comprehensive results, we separately compare the translation results for T1 $\to$ T2 and T2 $\to$ FLAIR, as some previous methods are only capable of single-modality transfer.
-These include both 2D and 3D generation methods, as shown in Tab.
-1.
-Additionally, we employed SwinUNTER for multi-modality translation comparisons.
-All evaluations were performed on 3D volumes; for the 2D methods, synthesized target images were stacked to form a 3D volume for comparison.
-Please refer to model training details to Appendix D.
+Figure 4: Generated samples on AUTOPET-II.
 Results.
 Tab.
-1 exhibits the transfer results of T1 $\to$ T2 and T2 $\to$ Flair.
+1 exhibits the transfer results in comparison between our and previous one-to-one modality transfer methods.
+Following previous of T1 $\to$ T2 and T2 $\to$ Flair.
 Our approach significantly surpasses previous 2D and 3D generation methods, including single- and multi-modality translation methods.
 Specifically, our approach exceeds current SOTA diffusion-based methods, such as 2D-based MD-Diff and 3D-based MS-SPADE.
-In terms of multi-modality translation, as Tab.
-2 shows, our approach performs better than MS-SPADE and SwinUNETR across all metrics under all settings.
-Moreover, for all settings, it can be seen that our method significantly improves the SSIM, indicating a better anatomy structure obtained by our approach.
+In terms of one-to-all multi-modality translation, as Tab.
+2 and additional results in the Appendix show, our approach with various volume sizes ($96$ and $192$) performs better than MS-SPADE and SwinUNETR across all metrics under all settings.
+Moreover, our method significantly improves the SSIM, indicating a better anatomy structure obtained by our approach.
 These results indicate that the $\mathbb{X}_{h}$ Hypothesis is plausible for homogeneous generalization, and our personalized approach is able to obtain its approximation.
-Analysis of $\mathbb{O}$.
-We show that using $\mathbb{O}$ for $\mathbb{X}_{h}$ mainly accomplishes the personalized knowledge of each sequence from MRI modalities.
-Those modalities are mainly focused on the physical anatomy.
-For the Flair modality in MRI, which mainly highlights the lesion but suppresses structures like bones, Fig.
-3 shows that without $\mathbb{O}$, the main difference between the generated images and ground truth (GT) images is the personalized structure.
-Prior $\mathbb{O}$ for $\mathbb{X}_{h}$ accomplishes and refines the personal level anatomical information, mitigating the gap between them with the GT, so it can be better transferred to other structural focusing modalities.
-Method Dice↑ Dice-↑ TPR↑ TNR↑ FNR↓ FPR↓
-From scratch
-nnUnet [21] 33.10 - - - - -
-UNETR [18] 10.81 23.14 80.65 3.64 19.35 96.36
-SwinUNETR [17] 43.45 62.60 90.32 62.73 9.68 37.27
-SwinUNETR with pre-train
-With pre-train [40] 44.06 57.79 89.25 73.64 10.75 26.36
-With ours 48.20 61.16 88.17 77.27 11.83 22.72
-Table 5: Segmentation results of PET and CT on AutoPET-II: Comparison between the previous method and ours. The best results are highlighted in blue.
-4.2 Tuning for missing modality segmentation
+Please refer to more analysis on $\mathbb{O}$ in Sec.
+G.1.
+Figure 5: T-SNE visualizations on embeddings of SwinUnetr and ours trained on BRATS23 for one-to-all modalities generation.
+More analysis on personalization.
+Fig.
+5 also shows that combining the prior and constraints results in more scattered embeddings at the individual level compared to the method without them.
+Tab.
+6, Tab.
+3, and the quantitative results in the paper further confirm the advantages of personalization.
+4.1.2 Fine-tuning stage: Missing modality segmentation
 Experimental settings.
 To validate the generalization ability of the pre-trained model, we fine-tune the model obtained from Sec.
-4.1 on the BRATS18 [32] from the Multimodal Brain Tumor Segmentation Challenge.
+4.1.1 on the BRATS18 [35] from the Multimodal Brain Tumor Segmentation Challenge.
 Similar to BRATS23, BRATS18 also consists of the same four structural modalities.
 We employ the Dice similarity coefficient (DICE) as the metric for evaluation.
-For a fair comparison, we follow data splits of [38] and reproduce the results of previous methods  on these splits by using their released code and following their original settings  [^2] Though we tried our best, it can be noticed some reproduced results are lowered than their reported results in their original paper.
+For a fair comparison, we follow data splits of [42] and reproduce the results of previous methods [14, 57, 46, 31, 42] on these splits by using their released code and following their original settings  [^2] Though we tried our best, it can be noticed some reproduced results are lowered than their reported results in their original paper.
 It should be clarified that our results also exceed those reported results.
@@
 .
-Additional experimental details can be seen in Appendix D.
+We also reproduce the results by using the SwinUNETR pre-training method [44], employing their official code to train the model on the same datasets with identical training epochs, learning rates, and other hyperparameters as ours.
+See additional experimental details in Appendix F.
 Results.
@@
 3 presents the segmentation results of our approach compared to previous methods.
-We also compute the standard deviation of DICE scores under various missing modality settings, which highlights the robustness of our model.
-Notably, our approach outperforms previous methods in most missing modality scenarios, particularly when the number of missing modalities is large.
-Our method shows significant improvements in enhancing tumor segmentation, especially when Missing Num = 3.
-Moreover, the reduced standard deviation of DICE scores under different missing modality settings indicates that our personalized approach consistently delivers superior segmentation results.
+We also compute the standard deviation of DICE scores under various missing modality settings, which highlights the robustness of models.
+Notably, our approach outperforms previous methods in most missing modality scenarios, particularly when the number of missing modalities is large (e.g., NM=3).
+Meanwhile, our approach also obtain a relatively low standard deviation for most settings, showing its robustness to various modalities missing scenario.
+Compared with SwinUNETR pre-training, our methods yield consistently better DICE with comparable standard deviation.
 This performance improvement stems from the enhanced generalization of our model, which is rooted in the learned ${\mathbb{X}_{h}}^{\prime}$.
-5 Heterogeneous generalization: PET and CT
-Given the differing imaging principles, the modality gap in heterogeneous generalization may be more pronounced than that in homogeneous generalization, making the former tasks more challenging.
-In this section, we evaluate our approach for heterogeneous generalization.
-5.1 Pre-training for modality transfer
+4.2 Heterogeneous generalization: PET and CT
+4.2.1 Pre-training stage: Modality transfer
 Experimental settings.
-We utilize the AutoPET-II dataset [15] from the Automated Lesion Segmentation in PET/CT challenge for pre-training.
+We utilize the AutoPET-II dataset from the Automated Lesion Segmentation in PET/CT challenge [17] for pre-training.
 The AutoPET-II dataset includes AC-PET and CT pairs, where the PET scans adopt FDG tracers, and their attenuation is corrected using the corresponding CT scans.
@@
 We use both contrastive loss and the decomposition constraint as our baseline.
-Please refer to training details in Appendix D.
+Please refer to training details in Appendix F.
 SSIM↑ PSNR↑
@@
 SwinUNETR 0.5853 0.9265 0.7559 23.5628 42.5495 33.0561
-Ours 0.6939 0.9516 0.8227 25.8498 46.4658 36.1578
+Ours (PUIR) 0.6939 0.9516 0.8227 25.8498 46.4658 36.1578
 SSIM↑ PSNR↑
@@
 SwinUNETR 0.4476 0.8705 0.6591 22.5086 41.3272 31.9179
-Ours 0.4744 0.8853 0.6798 22.7791 42.7687 32.7739
+Ours (PUIR) 0.4744 0.8853 0.6798 22.7791 42.7687 32.7739
 Table 5: Modality transfer results of NAC-PET to AC-PET and CT that tuned on HNSCC and evaluated on HNSCC validation set and NSCLC: Comparison between the previous method and ours for transfer between different modalities.
@@
 Figure 6: Modality transfer results of NAC-PET to AC-PET: Generated examples on the NSCLC dataset for NAC $\to$ AC across individuals.
-Figure 5: Generated samples on AUTOPET-II.
 Results.
@@
 4 and generated examples in Fig.
-5, incorporating $\mathbb{O}$ with different combinations of constraints improves generation quality across most metrics.
+4, incorporating $\mathbb{O}$ with different combinations of constraints improves generation quality across most metrics.
 Specifically, using the constraints without $\mathbb{O}$ does not guarantee improvements, as discussed in Sec.
-3.2.1.
+3.1.2.
 Ultimately, employing all constraints along with $\mathbb{O}$ yields the best average results across all translations, validating that our approach performs well in heterogeneous generalization settings.
 These results indicate that our method under the scope of personalization bridges the gap between structural and functional modalities.
-We validate the transferability of all these pre-train models in Sec.
-C.1, where additionally analysis are provided.
-5.2 Tuning for segmentation
+We validate the transferability of all these pre-trained models in Appendix E, where additional analyses are provided.
+4.2.2 Fine-tuning stage: Segmentation
 Experimental settings.
@@
 It is important to note that we employ the same training and testing splits as in Sec.
-5.1 to avoid data leakage.
+4.2.1 to avoid data leakage.
 Specifically, we adhere to the settings from the official challenge; DICE is calculated in the standard manner but is set to zero for false negatives and true negatives.
@@
 Notably, we also compare our approach with SwinUNETR using its pre-training strategy [44].
-Please refer to training details in Appendix D.
+Please refer to training details in Appendix F.
+From scratch Dice
+nnUnet [23] 33.1
+SwinUNETR [19] 43.5
+With 3D medical image SSL
+SwinUNETR [44] 44.1
+SwinMM [49] 44.2
+PCRL v2 [59] 41.9
+VoCo [51] 46.3
+Ours (PUIR) 48.2
+Table 6: Comparison between different SOTA medical 3D SSL methods on AUTOPET segmentation.
 Results.
-The full modality segmentation results are exhibited in Tab.
-5.
-The results indicate that with proper model architecture, such as SwinUNETR, using both two modalities usually outperforms solely using PET.
-It can be observed that models using our pre-train improve the results across all metrics.
-Typically, SwinUNETR using our pre-train significantly exceeds it without our pre-trained model, indicating the personalized invariant learned by our pre-train generalizes to the downstream well and can boost the downstream tasks.
-Moreover, using our proposed components with the pre-train leads to the best DICE and DICE-.
-This validates that using the prior further emphasizes the personalized invariant, which yields the most segmentation improvements.
-6 Special case: A complex scenario
-We introduce a more complex scenario, in which the pre-trained model for heterogeneous generalization settings is tuned downstream that span both heterogeneous and homogeneous generalization.
+The segmentation results are presented in Tab.
+6, with additional details provided in Appendix Tab.
+8.
+For a fair comparison, all other SSL methods were reproduced using the same pre-training and fine-tuning datasets as those used in our approach.
+Our approach significantly improves the DICE results in comparison to other SSL methods since we specifically address the cross-modality generalization.
+The performance gains on heterogeneous modalities also enhance the significance of our $\mathbb{X}_{h}$ hypothesis for heterogeneous generalization.
+4.3 Fine-tuning special case: A complex scenario
+We introduce a more complex scenario in which the pre-trained model for heterogeneous generalization settings is tuned downstream to span both heterogeneous and homogeneous generalization.
 Experimental settings.
 The pre-train model we adopted is from Sec.
-5 that trained on AC-PET and CT.
+4.2 that trained on AC-PET and CT.
 Specifically, we tune the model by using the Head and Neck Squamous Cell Carcinoma (HNCSS) dataset [18] as the training set and the Non-Small Cell Lung Cancer (NSCLC) dataset as the testing set.
@@
 Note here the training and testing data in the downstream task come from different domains.
-See training details in Appendix D.
+See training details in Appendix F.
 Results.
@@
 We demonstrate that learning $\mathbb{X}_{h}$ is both feasible and highly beneficial for enhancing generalization in medical tasks.
-By discussing limitations, future directions, and social impact in Appendices B and C, our study may point to a promising path for achieving medical generalization through personalization in complex settings.
-Appendix A Related work
+Our method also correlates to personalized medicine, a transformative framework for $21^{st}$ century healthcare, tailoring medical treatments to each patient’s unique characteristics [50, 26, 5].
+By discussing limitations, future directions, and social impact in Appendices C and D, our study may point to a promising path for achieving medical generalization through personalization in complex multi-modality medical analysis.
+Acknowledgments
+This work was supported by the National Natural Science Foundation of China (Grant No.
+92370119, No.
+62376113, No.
+62206225, No.
+62436009, No.
+62276258, No.
+82394432, and No.92249302), the Shanghai Municipal Science and Technology Major Project (Grant No.
+2023SHZDZX02), and XJTLU Funding REF-22-01-002.
+The computations in this research were performed using the CFFF platform of Fudan University.
+Supplementary Material
+Appendix A Brief descriptions of different modalities
+Magnetic Resonance Imaging (MRI) scans [58]: use strong magnetic fields and radiofrequency currents yielding distinct sequences.
+Typically, MRI has different modalities, include T1, T2, T1ce and Flair.
+Computed Tomography (CT) scans [36, 56] employ X-rays to measure its attenuation.
+Positron Emission Tomography (PET) scans are expensive functional imaging scans that employ radiotracers emitting gamma rays to visualize and measure metabolic processes.
+Thus, PET scans have a large percentage of background areas.
+Appendix B Related work
 Medical generalization tasks.
@@
 Our work reveals that our pre-train model with basic segmentation tuning exceeds these approaches.
-Appendix B Limitations, challenges, and future work
+Appendix C Limitations, challenges, and future work
 To enhance the validation of our approach, we adhere to commonly used settings during the tuning stage.
@@
 Additionally, we advocate for the availability of more open-source multi-modal medical datasets, particularly for functional modalities, as these are not widely accessible to researchers.
-Appendix C Social impact
+Appendix D Social impact
 This work presents an approach to tackle multi-modality generalization through personalization.
 We hope our work can encourage the community to work towards practical, personalized medical models with border generalization ability.
-C.1 Downstream segmentation ablation study
+Appendix E Downstream segmentation ablation study
 Table 7: Ablation study - Segmentation results of using different pre-train models on AutoPET-II: Comparison between the pre-train models with different settings and ours.
@@
 48.20 61.16 88.17 77.27 11.83 22.72
-We demonstrate the effectiveness of our proposed components and discuss the process of learning an anatomy-invariant representation.
+The effectiveness of our proposed components is demonstrated alongside an exploration of the methodology employed to develop an individual-invariant representation.
 Experimental results for downstream segmentation tasks and visualizations of the pre-trained models are presented in Tab.
@@
 Consistent with Sec.
-3.2.1, the results indicate that using different constraints alone may not guarantee improvements; however, incorporating all constraints along with $\mathbb{O}$ results in the best outcomes.
-This validates the plausibility of the $\mathbb{X}_{h}$ Hypothesis and demonstrates that achieving good approximation of it significantly enhances generalization.
+3.1.2, the results indicate that using different constraints alone may not guarantee improvements; however, incorporating all constraints along with $\mathbb{O}$ results in the best outcomes.
+This validates the plausibility of the $\mathbb{X}_{h}$ Hypothesis and demonstrates that achieving a good approximation of it significantly enhances generalization.
 Using prior $\mathbb{O}$ with decomposition constraint improves the model performance for different settings.
@@
 Combined with the improvements from modality transfer results in Tab.
-4, it suggests that $\mathbb{O}$ helps with better obtaining anatomical structure.
+4, it suggests that $\mathbb{O}$ helps in better obtaining anatomical structure.
 The invariance and equivariance constraints can not be applied to the same feature.
@@
 As shown in task 3, without $\mathbb{O}$, invariance and equivariance constraints are applied to the latent feature simultaneously, leading to a significant performance drop.
-In comparison, apply equivariance constraint before using $\mathbb{O}$ and applying invariance constraint after using $\mathbb{O}$ yields the best results.
-This is because the geometrical transformation contained in $z^{i}_{h}$ needs to be accomplished by fetching other possible geometrical transformation information form $\mathbb{O}$ and then fused to be invariant.
-Appendix D Experimental details
+In comparison, applying equivariance constraint before using $\mathbb{O}$ and applying the invariance constraint after using $\mathbb{O}$ yields the best results.
+This is because the geometrical transformation contained in $z^{i}_{h}$ needs to be accomplished by fetching other possible geometrical transformation information from $\mathbb{O}$ and then fusing it to be invariant.
+Appendix F Experimental details
 The model and data loaders are built by using MONAI https://docs.monai.io/en/stable/index.html.
@@
 We present some key implementations below.
-D.1 Overall training procedure
-We provide a pseudo-code for our approach.
+F.1 Overall training procedure
+A pseudo-code is provided for our approach.
 The loss calculation for Pre-training procedure is simplified as Algorithm 1 and Downstream tuning as Algorithm 2.
@@
 Algorithm 2 Calculate losses during one step for fine-tuning
-D.2 Homogeneous generalization: structural modalities in MRI
-D.2.1 Pre-training and Modality transfer.
+F.2 Homogeneous generalization: structural modalities in MRI
+F.2.1 Pre-training and Modality transfer.
 Experimental settings.
-We use four A100 GPUs for training.
+Four A100 GPUs are employed for training.
 The learning rate we used for the modality transfer is set to $0.0002$, and the training epoch is set to $1000$.
@@
 The training loss follows the $\mathcal{L}_{pre}$, whose calculation details during the training phase can be seen in Algorithm 1.
-D.2.2 Missing modality segmentation.
-We use four A100 GPUs for tuning.
+F.2.2 Missing modality segmentation.
+Four A100 GPUs are employed for tuning.
 The learning rate we used for the modality transfer is set to $0.0002$, and the training epoch is set to $1000$.
@@
 For the fine-tuning, we employ Dice loss, the weighted cross-entropy loss that is adopted by [42], and the additional $\mathcal{L}_{inv}$.
-D.3 Heterogeneous generalization: PET and CT modalities
-D.3.1 Modality transfer
+F.3 Heterogeneous generalization: PET and CT modalities
+F.3.1 Modality transfer
 All models are trained using A100 GPUs. Training details. All models are trained under the same situations, using the same data pre-processing transforms.
-D.3.2 Downstream segmentation
+F.3.2 Downstream segmentation
 Training details.
@@
 Moreover, we also compare the original architecture of SwinUNETR using our pre-trained weights with fully using our architecture and our weights for fine-tuning.
-D.4 Special case: Tuning from heterogeneous to homogeneous generalization with domain gap
+F.4 Fine-tuning special case: Tuning from heterogeneous to homogeneous generalization with domain gap
 Training details.
@@
 Additionally, we reproduced the results of UNETR and SwinUNETR for comparison, ensuring that the same loss functions were applied across models.
-Appendix E More results
-E.1 Modality transfer results on BRATS22:
+Appendix G More results and analysis
+G.1 More analysis on learnable biological prior
+Analysis of $\mathbb{O}$.
+We show that using $\mathbb{O}$ for $\mathbb{X}_{h}$ mainly accomplishes the personalized knowledge of each sequence from MRI modalities.
+Those modalities are mainly focused on the physical anatomy.
+For the Flair modality in MRI, which mainly highlights the lesion but suppresses structures like bones, Fig.
+8 shows that without $\mathbb{O}$, the main difference between the generated images and ground truth (GT) images is the personalized structure.
+Prior $\mathbb{O}$ for $\mathbb{X}_{h}$ accomplishes and refines the personal level anatomical information, mitigating the gap between them with the GT, so it can be better transferred to other structural focusing modalities.
+Figure 7: Visualizations of generated modalities with T1 as input of our method, which allows the capturing of subtle structures.
+Figure 8: Visualization of the efficacy of prior $\mathbb{O}$.
+Displayed are the generated modalities on the input Flair modality of a testing sample on the BTATS21 dataset.
+Columns show: the generated images of the model (1) without prior $\mathbb{O}$ and (2) with prior $\mathbb{O}$ are aligned with (3) the GT images.
+Typically, the differences between without and with prior $\mathbb{O}$ (the (2)-(1) column) are visualized to compare with the differences between without $\mathbb{O}$ and GT (the (3)-(1) column).
+Red and blue refer to the positive (accomplishment) and negative (refinement) values of the differences, respectively.
+G.2 Segmentation results on AutoPET-II.
+Detailed metrics results of AutoPET-II are presented in Tab.
+8.
+The results indicate that with proper model architecture, such as SwinUNETR, using both two modalities usually outperforms solely using PET.
+It can be observed that models using our pre-train improve the results across all metrics.
+Typically, SwinUNETR using our pre-train significantly exceeds it without our pre-trained model, indicating the personalized invariant learned by our pre-train generalizes to the downstream well and can boost the downstream tasks.
+Moreover, using our proposed components with the pre-train leads to the best DICE and DICE-.
+This validates that using the prior further emphasizes the personalized invariant, which yields the most segmentation improvements.
+Method Dice↑ Dice-↑ TPR↑ TNR↑ FNR↓ FPR↓
+From scratch
+nnUnet [23] 33.10 - - - - -
+SwinUNETR [19] 43.45 62.60 90.32 62.73 9.68 37.27
+SwinUNETR with different pre-train
+With pre-train in [44] 44.06 57.79 89.25 73.64 10.75 26.36
+With ours 48.20 61.16 88.17 77.27 11.83 22.72
+Table 8: Segmentation results of PET and CT on AutoPET-II: Comparison between the previous method and ours. The best results are highlighted in blue.
+G.3 Modality transfer results on BRATS22.
 Tab.
@@
 Generated examples are presented in Figs.
-6, 7 and 8.
-Table 8: Modality transfer results of MRI on BRATS23: Comparison between previous methods and our method. The best results are highlighted in blue.
+10, 11 and 12.
 Task T1→T2 T2 → Flair
@@
 3D Ours 30.756 ± 1.950 0.065 ± 0.034 0.944 ± 0.031 32.224 ± 2.518 0.046 ± 0.029 0.941 ± 0.025
-Table 9: Modality transfer results of MRI on BRATS23: The averaged results with standard derivations of metrics between all modalities.
+Table 9: Modality transfer results of MRI on BRATS23: Comparison between previous methods and our method. The best results are highlighted in blue.
 Target T1 T1ce T2 Flair
@@
 Flair Std. 2.399 0.020 0.022 1.634 0.035 0.021 1.951 0.035 0.034 2.441 0.008 0.013
-E.2 Missing modality segmentation results on BRATS18:
+Table 10: Modality transfer results of MRI on BRATS23: The averaged results with standard derivations of metrics between all modalities.
+G.4 Missing modality segmentation results on BRATS18.
 We provide detailed segmentation results on BRATS18 as Tab. 11.
-Table 10: Missing modality segmentation results of MRI on BRATS18: Num denotes the number of missing modalities for different settings.
-The used modalities are highlighted with gray boxes and the missing ones remain as blank.
-The results of each setting are presented accordingly.
+G.5 Comparison with non-personalized methods.
+We provide further visual evidence in Fig.
+9 for an in-depth analysis.
+Fig.
+9 shows that applying our constraints aligns CT and PET representations more closely, indicating smaller $d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{M}|\mathcal{H}^{\mathcal{S}},\mathcal{M}|\mathcal{H}^{\mathcal{U}})$ in Paper Eq.
+(10), yielding tightened bounds and better downstream performance (Dice:48.20), in comparison to the non-personalized baseline (Dice:43.5).
+This further supports our theoretical analysis.
+Figure 9: TSNE of latent features of PET and CT images obtained under different applied constraints. Downstream performances are noted.
 Missing Num =3 =2 =1 =0
@@
 Whole tumour Ours 89.23 81.73 82.26 87.45 89.74 89.03 88.00 81.92 89.72 86.27 89.12 90.5 91.25 87.1 91.19
+Table 11: Missing modality segmentation results of MRI on BRATS18: Num denotes the number of missing modalities for different settings.
+The used modalities are highlighted with gray boxes and the missing ones remain as blank.
+The results of each setting are presented accordingly.
 Figure 10: Generated images of our proposed method: slices across ventricles.
 Figure 11: Generated images of our proposed method: slices across cerebral sulcus.
-Figure 8: Generated images of our proposed method: slices across cerebellar hemisphere. Our method can generate defined cerebellar folia.
+Figure 12: Generated images of our proposed method: slices across the cerebellar hemisphere. Our method can generate defined cerebellar folia.